### v1

Abstract
Generative AI is a directional technology: it excels at some task combinations and performs poorly at others. Knowledge work is also directional and endogenous: workers can satisfy their job requirements with different combinations of tasks. Studying AI adoption by knowledge workers hence requires comparing two vectors. We develop a high-dimensional model of task choice and technology adoption, with otherwise standard neoclassical assumptions. AI is adopted when its direction is aligned with what the worker values at the margin—the worker’s shadow prices, rather than with what the worker actually does—their activity vector. This yields a cone of adoption that widens as AI capability grows; near the entry threshold, small improvements in capability translate into large expansions in the set of adopted directions. Adoption also has a structured intensive margin: a tool can be worth using but not worth using all the time, generating a region of stable hybrid production between an entry threshold and an all-in threshold. We also show how to derive shadow prices as explicit functions of observable skill and requirement vectors. The framework explains rapid adoption in aligned occupations, heterogeneous adoption elsewhere, and weak correlation with one-dimensional skill measures: the key heterogeneity is directional alignment, not skill level.
1  Introduction
Generative artificial intelligence (GenAI) based on large language models (Vaswani et al., 2017) shares features of general-purpose technologies (Eloundou et al., 2023) but differs from earlier automation in two ways. First, individual workers can adopt GenAI quickly in knowledge work: the interface is general-purpose and the main inputs are time and attention rather than specialized capital. Adoption can therefore spread through informal and unregulated channels inside organizations. [^1] On worker-driven adoption (”shadow IT”), see Klotz et al. (2019); Raković et al. (2020); for rapid early diffusion, see UBS (2023). Second, capabilities improve quickly but unevenly across tasks. [^2] For software engineering benchmarks, see SWE-bench (Jimenez et al., 2024). The unevenness is described as a ”jagged technological frontier” by Dell’Acqua et al. (2023) and quantified more broadly by Hendrycks et al. (2025). Taken together, these features make it harder to predict who adopts AI and who gains from it.
Measured productivity effects are mixed and do not map cleanly to any single notion of skill. In some settings, usage is already pervasive and mediated by tools that translate intent into working output; in others, adoption is slower and gains depend on complementary investments by firms. [^3] Evidence on productivity effects is mixed across settings and designs; see, e.g., Dell’Acqua et al. (2023); Benzell and Myers (2026) and firm-level evidence on complementarity in Aldasoro et al. (2026). Existing frameworks often compress heterogeneity into task or worker groups or a scalar exposure index, which can obscure the direction-by-direction nature of adoption and motivate a multidimensional model.
To address this, we develop a model where both jobs and AI systems are explicitly multidimensional. A job is a bundle of tasks (Autor, 2013, 2025), and a worker chooses her task mix to maximize output subject to a time or attention budget. Skills shape the feasible set of task combinations, while job requirements shape the value of tasks in production. We model AI as a directional technology that produces tasks in fixed proportions. Using AI requires supervision—time the worker cannot spend on her own production. Because supervision draws on the same scarce budget, adoption is a team-production problem: the combined feasible set is a weighted sum of the worker’s own capabilities and those of AI. When their capabilities point in different directions, the joint production set expands beyond what either achieves in isolation. [^4] The logic extends to AI that can flexibly combine a small number of task directions in a high-dimensional space; see Kaushik et al. (2025). The key analytical step is to evaluate a technology direction at the worker’s marginal valuation: at the optimum, shadow prices convert any task direction into a payoff-relevant scalar.
The model yields a simple organizing prediction: adoption is governed by alignment between the AI direction and what the worker values at the margin. AI is adopted when its direction is sufficiently aligned with the worker’s shadow prices, which are in general different from the worker’s observed activity mix.
We make six contributions.
  1. 1.
    Explain nonlinear adoption without frictions. The model generates a sharp capability threshold for adoption even in a frictionless neoclassical setting. Near this threshold, small improvements in AI capability produce large expansions in the set of adopted tasks. This provides a clean economic explanation for why usage can suddenly take off without invoking hype, network effects, or fixed costs.
  2. 2.
    Separate absolute from comparative advantage. Technological superiority is necessary but not sufficient: AI is used only when its strengths align with what is scarce at the margin in a given worker–job match. This explains why powerful tools are sometimes ignored and why adoption varies widely across firms and roles.
  3. 3.
    Reframe AI as team production. Because AI requires supervision and consumes time and attention, it competes for the same scarce resource as human work. Adoption is therefore a collaboration decision evaluated at shadow prices, not a simple capital–labor substitution story. This shifts the discussion from replacement to directed co-production.
  4. 4.
    Introduce a structured intensive margin. There are two adoption thresholds in our model: an entry threshold and an all-in threshold. Between them, optimal AI use is partial, generating stable hybrid production rather than immediate automation.
  5. 5.
    Establish the worker–job match as the unit of analysis. Adoption depends on the interaction between worker skills and job requirements, not on workers or occupations in isolation. Shadow prices are match-specific, so asking which jobs or which workers are at risk is incomplete without specifying the pairing.
  6. 6.
    Break the scalar exposure paradigm. Adoption cannot be summarized by a one-dimensional skill or exposure index. The relevant sufficient statistic is a vector of shadow prices reflecting multidimensional alignment. As a result, common exposure rankings may systematically misclassify who adopts and who benefits.
These results imply that scalar exposure rankings are generally not sufficient statistics for adoption.
To make the mechanism transparent, our method is intentionally neoclassical. We study a worker doing a given job and abstract from governance and principal-agent frictions, so the worker maximizes output. We characterize autarky task choice using a trade-style duality argument (Dixit and Norman, 1980), which yields shadow (autarky) prices even though tasks are not traded on any market. We then introduce AI as a directional technology that competes for the same scarce budget and derive adoption and intensity conditions in terms of capability and alignment.
For interpretation and measurement, we provide a CES-CET parametric specialization that expresses shadow prices as explicit functions of skill and requirement vectors. The key parameters are the elasticity of substitution across tasks in production ($\sigma$) and the elasticity of transformation in resource use ($\gamma$). When $\gamma\to 0$, workers are rigid specialists and skills alone determine autarky prices; when $\gamma\to\infty$, workers are perfectly flexible and job requirements alone matter. Symmetric logic applies to $\sigma$: low values make tasks strong complements, high values make them close substitutes. In general, both worker skills and job requirements shape autarky prices, consistent with within-job task reallocation (Spitz-Oener, 2006).
The framework also has limitations that suggest where the economics of GenAI must ultimately go. By design, we hold the worker-AI match fixed and treat adoption as an individual optimization problem. We do not model sorting—which tools a worker selects and which workers a tool is designed for—or how organizations structure human-AI teams. In practice, firms set policies, provide complementary investments, and may restrict or encourage usage (Aldasoro et al., 2026). We also treat the technology as exogenous and focus on the worker side of adoption rather than on how training data and market incentives shape the direction of AI capabilities. These abstractions isolate a basic mechanism—directional alignment at shadow prices—that can be embedded in richer models of organizational choice, market structure, and dynamic capability change.
The remainder of the paper proceeds as follows. Section 2 develops the model of task choice and directional technology adoption, derives the adoption condition and the cone of adoption, and characterizes the intensive margin. Section 3 provides a CES-CET parametric specification that yields closed-form expressions for shadow prices and adoption thresholds. Section 4 positions the framework relative to prior task-based models and discusses interpretation, limitations, and connections to the emerging evidence. The Appendix collects proofs and the parametric derivations.
2  A Model of Digital Technology Adoption
We study a worker doing a given job. We do not model labor market, governance, or principal-agent frictions; consequently, the worker maximizes output. This is the outcome of a richer model in which both employer and worker can observe and contract on task execution. We take the worker-job match as fixed and reserve the study of sorting to future work.
The worker chooses how much to do of each of $N$ tasks. Denote the task quantities by the vector $x\in\mathbb{R}_{+}^{N}$. The set of task bundles the worker can complete in one workday (or other unit of scarce resource) is given by the set $X_{A}$, characterized by its resource use function $g$,
$X_{A}=\{x:g(x)\leq B\}.$
Here $B$ is the amount of the scarce resource (“budget”) the worker has available. We interpret the budget as time or attention, not money. This is a key feature of digital technology adoption: the binding constraint is typically the worker’s cognitive capacity and hours, not financial cost. The set $X_{A}$ is called the Production Possibilities Set (PPS).
Assumption 2.1 (Resource use function).
The resource use function $g:\mathbb{R}^{N}_{+}\to\mathbb{R}_{+}$ is (i) convex, (ii) homogeneous of degree one, (iii) increasing in each component, and (iv) $g(0)=0$. These properties make the feasible set $X_{A}=\{x:g(x)\leq B\}$ convex, scalable with the budget, and closed under free disposal. We do not require $g$ to be differentiable or strictly convex.
Assumption 2.2 (Production function).
The production function $F:\mathbb{R}^{N}_{+}\to\mathbb{R}_{+}$ is (i) homogeneous of degree one (constant returns to scale), (ii) strictly quasi-concave (upper contour sets $\{x:F(x)\geq c\}$ are strictly convex), and (iii) twice continuously differentiable on $\mathbb{R}^{N}_{++}$ with $\nabla F(x)\neq 0$ for all $x\in\mathbb{R}^{N}_{++}$. Strict quasi-concavity ensures that the optimum, when interior, is unique and the supporting hyperplane is pinned down by $\nabla F$.
Definition 2.1 (Task allocation problem).
Given a resource use function $g$, a production function $F$, and a budget $B>0$, the worker’s task allocation problem is
$\max_{x\in\mathbb{R}^{N}_{+}}F(x)\qquad\text{s.t.}\qquad g(x)\leq B.$
The worker chooses the task bundle that maximizes output subject to the constraint that total resource use does not exceed the budget.
Proposition 2.1 (Existence).
Under Assumptions 2.1 and 2.2, a solution to the task allocation problem (Definition 2.1) exists.
Proof.
The feasible set $X_{A}$ is closed (preimage of $(-\infty,B]$ under continuous $g$) and bounded (homogeneity and monotonicity of $g$ imply $g(x)\to\infty$ as $\|x\|\to\infty$). The objective $F$ is continuous ($C^{2}$ implies continuity). Existence follows from Weierstrass’s theorem. ∎
We denote any solution by $x_{A}$.
Definition 2.2 (Shadow prices).
A shadow price vector $p\in\mathbb{R}^{N}_{+}$ is the outward normal to a hyperplane supporting the Production Possibility Set at the optimum. We normalize all price vectors to have unit Euclidean length, $\|p\|_{2}=1$.
Proposition 2.2 (Autarky price characterization).
A task bundle $x_{A}$ solves the task allocation problem if and only if there exists an autarky price vector $p_{A}$ such that (i) $x_{A}$ maximizes revenue $R=p_{A}^{\prime}x$ subject to $g(x)\leq B$, and (ii) $x_{A}$ maximizes output $F(x)$ subject to $p_{A}^{\prime}x\leq R$. Moreover, if $x_{A}\in\mathbb{R}^{N}_{++}$, then $p_{A}$ is unique.
Proof.
The first part follows from a standard tangency argument: at the optimum, a supporting hyperplane separates the feasible set from higher-output bundles, and its normal vector can be interpreted as implicit (autarky) task prices. When $x_{A}\in\mathbb{R}^{N}_{++}$, strict quasi-concavity of $F$ (Assumption 2.2) implies that $\nabla F(x_{A})$ uniquely pins down the supporting hyperplane, making $p_{A}$ unique even if $g$ is not differentiable at $x_{A}$. ∎
Figure 1 shows optimal task allocation with smooth $g$ and $F$ functions. The worker chooses the task bundle $x_{A}$ that maximizes output $F(x)$ subject to the Production Possibility Set (PPS). The autarky price vector $p_{A}$ is normal to the supporting hyperplane at the optimum. The boundary of the PPS is called the Production Possibility Frontier (PPF).
[…]
Figure 1: Task allocation in autarky
For notational convenience, we introduce the unit revenue function as
$\varrho(p)=\frac{1}{B}\max_{x:g(x)\leq B}p^{\prime}x=\max_{x:g(x)\leq 1}p^{\prime}x$
that shows, for every possible set of prices, the maximum “revenue” the worker can achieve using only human activities for one unit of time (or resource). [^5] Geometrically, this function represents a directional diameter of the PPS $X_{A}$. While the extent of $X_{A}$ is limited in each direction $x$ by $g(x)\leq B$, this function converts the directional extent into an inner product. Clearly, $\varrho(p)$ is homogeneous of degree one and convex.
2.1  Technology expands the production possibilities set
We model AI as a directional technology capable of executing tasks in fixed proportions. Concretely, AI produces tasks in a fixed mix $t$, scaled up by an intensity $\lambda\geq 0$. We normalize $\|t\|=1$ so that $t$ only represents a direction, not a length. Elements are $t_{i}\geq 0$ but may be zero in any particular direction. For example, AI may not be able to help in manual tasks.
We first discuss the properties and the effects of a single technology. We generalize the technology adoption problem to multiple directional technologies in Section 2.4.
To use AI, the worker has to use some of the same resources that constrain them in their own work. For example, instructing AI agents and reviewing their work takes time, and this limits the amount of AI that can be used in a workday. [^6] Time as a binding constraint is central to the economics of digital goods. Goodman et al. (2026) leverage Becker’s time allocation theory to study consumer demand for time-intensive platforms like Facebook and Instagram, finding that time shares and time costs significantly influence substitution patterns across online activities. We assume AI is competing for the same resource budget $B$, so if the worker spends a fraction $\lambda$ of their budget working with AI, only $(1-\lambda)B$ is left for doing other tasks. [^7] For simplicity, we assume there is a linear transformation between resources spent on AI supervision and resources spent on human work. This is eminently true if the resource is time: one hour lost on an activity is one hour gained on another. A more general formulation could model the production possibilities set as constrained by several resources, such as time, attention, mental energy, and AI may not use them in the same proportion as human tasks. We conjecture our results would survive as long as AI is sufficiently substitutable to human tasks in the resource constraint. We denote the capability of technology with $\chi\geq 0$, which acts as a productivity shifter. A unit of technology $\|t\|$ costs $1/\chi$ units of the resource $B$. Continuing with the AI example, more capable AI tools can get more done per hour of human supervision.
Definition 2.3 (Technology).
A technology is a pair $(t,\chi)$. The vector $t\in\mathbb{R}^{N}_{+}$ with $\|t\|_{2}=1$ is the direction (which tasks the technology performs, and in what proportions). The scalar $\chi\geq 0$ is the capability (how many units of tasks the technology produces per unit of resource spent supervising it).
Allocating $r$ units of resource to technology yields $r\chi t$ units of tasks. When the worker allocates a fraction $\lambda\in[0,1]$ of their budget $B$ to technology, they obtain $\lambda B\chi t$ from technology and have $(1-\lambda)B$ resources left for human work.
Definition 2.4 (Technology Adoption Problem).
Given technology $(t,\chi)$ and budget $B$, the worker chooses technology intensity $\lambda\in[0,1]$ and human task allocation $x_{H}$ to maximize output $F(x_{H}+\lambda B\chi t)$ subject to $g(x_{H})\leq(1-\lambda)B$.
The Production Possibility Set with technology is the set of all task bundles the worker can achieve by combining human work with (possibly partial) use of technology, allowing free disposal of any tasks:
$X_{T}=\left\{z\in\mathbb{R}^{N}_{+}:z\leq x_{H}+\lambda B\chi t,\;g(x_{H})\leq(1-\lambda)B,\;\lambda\in[0,1]\right\}.$
The inequality $z\leq x_{H}+\lambda B\chi t$ (componentwise) reflects free disposal: the worker can always discard output in any task. The autarky PPS $X_{A}$ already satisfies free disposal (Assumption 2.1), but the technology contribution $\lambda B\chi t$ is a single point in task space. Without the free disposal clause, $X_{T}$ would not be downward-closed in the technology direction.
As can already be seen from the definition, technology expands the production possibilities set into a convex hull, as we state formally below. Figure 2 illustrates the geometry. The worker allocates a fraction $\lambda$ of the budget to using technology (vector $y$) and a fraction $1-\lambda$ to their own work. The optimal allocation lies on the tangent line from $y$ to the Production Possibility Set. Point $a$ represents own work, and the line segment from $a$ to $b$ represents the contribution of technology.
Proposition 2.3 (Technology expands PPS to convex hull).
The Production Possibility Set with technology adoption equals the convex hull of the autarky PPS and the technology point, closed under free disposal:
$X_{T}=\text{conv}(X_{A}\cup\{B\chi t\})-\mathbb{R}^{N}_{+},$
where the subtraction denotes the Minkowski difference (i.e., for any $z$ in the convex hull, all $z^{\prime}\leq z$ with $z^{\prime}\geq 0$ are also in $X_{T}$). Since $X_{A}$ is already downward-closed, this simplifies to $X_{T}=\text{conv}(X_{A}\cup\{B\chi t\})$ whenever the technology point $B\chi t$ is dominated componentwise by some element of $X_{A}$, and otherwise augments the convex hull by the free-disposal region below the technology point.
Proof.
Any $z\in X_{T}$ satisfies $z\leq(1-\lambda)\tilde{x}+\lambda(B\chi t)$ for some $\tilde{x}=x_{H}/(1-\lambda)\in X_{A}$ (using homogeneity of $g$), hence $z$ lies in the downward closure of the convex hull. Conversely, take any $w=(1-\mu)x+\mu(B\chi t)$ with $x\in X_{A}$, $\mu\in[0,1]$, and any $z\leq w$, $z\geq 0$. Set $\lambda=\mu$ and $x_{H}=(1-\mu)x$. Then $g(x_{H})=(1-\mu)g(x)\leq(1-\mu)B$ and $z\leq x_{H}+\lambda B\chi t$, so $z\in X_{T}$. ∎
[…]
Figure 2: Technology adoption as a convex combination
The new PPS is larger than the previous one whenever the technology vector “pierces” the old PPS, that is, technology has an absolute advantage over the worker in some combination of human tasks. When the entire technology vector is inside the worker’s PPS, they do not gain anything from it, because the resource spent on working with the technology is less productive than doing the same combination of tasks manually.
Let $x_{t}$ denote the combination of tasks that the worker could produce on their own in the exact direction of technology.
Definition 2.5 (Technology-direction bundle).
For technology direction $t$, define $x_{t}$ as the point on the PPF in direction $t$:
$x_{t}=\frac{B}{g(t)}t.$
This follows from homogeneity: $g(\epsilon t)=\epsilon g(t)$, so the maximum $\epsilon$ satisfying $g(\epsilon t)\leq B$ is $\epsilon^{*}=B/g(t)$.
Definition 2.6 (Absolute advantage).
Technology $(t,\chi)$ has absolute advantage over a worker with resource use function $g$ if
$\chi g(t)>1.$
Equivalently, $\|B\chi t\|>\|x_{t}\|$: technology can produce more in direction $t$ than the worker. Since $\|x_{t}\|=B/g(t)$ and $\|B\chi t\|=B\chi$, the condition reduces to $\chi>1/g(t)$. Note that absolute advantage is independent of budget $B$. Absolute advantage can arise either because technology is very capable ($\chi$ large) or because the worker is particularly inefficient in direction $t$ (large $g(t)$, meaning high resource use per unit of tasks in that direction).
Proposition 2.4 (Absolute advantage necessary for adoption).
If technology $(t,\chi)$ does not have absolute advantage ($\chi g(t)\leq 1$), then $X_{T}=X_{A}$: technology does not expand the PPS.
Proof.
Without absolute advantage, $B\chi\leq B/g(t)$, so $g(B\chi t)=B\chi g(t)\leq B$. Thus the technology point $y=B\chi t\in X_{A}$. Since $X_{A}$ is convex, $\text{conv}(X_{A}\cup\{y\})=X_{A}$. ∎
This proposition is illustrated in Figure 3. When the technology vector $y=B\chi t$ does not extend beyond the worker’s PPS, the worker does not adopt technology. The supporting hyperplane at the tangency point shows that the worker can achieve higher output using their own skills alone. [^8] The figures normalize $B=1$, so the technology vector is shown as $y=\chi t$.
[…]
Figure 3: Technology inside the Production Possibility Set ($\chi g(t)<1$)
Absolute advantage is necessary but not sufficient for technology adoption. A counterexample is illustrated in Figure 4. The technology already has an absolute advantage, so the worker could use it to expand their PPS. In their current job, however, their activities are directionally far from the technology, so it is not worth adopting it. [^9] Geometrically, until the technology vector pierces the budget set given by the hyperplane separate the PPS and the production function upper contour set, its convex hull with the PPS is also strictly within the budget set, extending no revenue increasing opportunities to the worker.
[…]
Figure 4: Technology has an absolute but not comparative advantage
2.2  The cone of technology adoption
Absolute advantage is necessary but not sufficient for adoption. The worker must also find technology directionally useful given their job requirements. We now characterize when adoption occurs.
Recall that $\varrho(p)$ is the unit revenue function, giving the implicit revenue the worker can achieve per unit of their resource when shadow prices are $p$. The “autarky revenue” can then be written as
$p_{A}^{\prime}x_{A}=B\varrho(p_{A}).$
For technology to be useful for the worker, it has to bring more implicit revenue.
Definition 2.7 (Alignment angle).
The alignment angle $\phi$ between autarky prices $p_{A}$ and technology direction $t$ is
$\phi=\arccos(p_{A}^{\prime}t)\in[0,\pi],$
where both $p_{A}$ and $t$ are unit vectors.
The worker is indifferent to adopting technology when the technology point $y=B\chi t$ brings exactly the same implicit revenue as the autarky activity vector. This happens when $y$ lies exactly on the supporting hyperplane at the autarky allocation, that is, when $p_{A}^{\prime}y=p_{A}^{\prime}x_{A}=B\varrho(p_{A})$. This pins down a threshold capability for each technology direction such that the worker adopts that technology if and only if it is more capable than the threshold. Larger capability always leads to adoption: if the worker adopts at $\chi$, the output-maximizing bundle under $\chi$ remains feasible under any $\chi^{\prime}>\chi$, so the worker continues to adopt.
Definition 2.8 (Adoption threshold).
For technology direction $t$ and shadow price $p$ with $p^{\prime}t>0$, define the adoption threshold
$\underline{\chi}(t,p)=\frac{\varrho(p)}{p^{\prime}t}.$
The adoption threshold is homogeneous of degree zero in prices and equals infinity when $p^{\prime}t=0$. Because $p$ and $t$ are unit vectors, their inner product $p^{\prime}t=\cos\phi$ is the cosine of the alignment angle.
The next theorem characterizes the adoption decision for a particular technology direction.
Theorem 2.5 (Technology adoption condition).
A worker with budget $B$, autarky prices $p_{A}$, and autarky allocation $x_{A}$ adopts technology $(t,\chi)$ if and only if
$B\chi\cdot p_{A}^{\prime}t>p_{A}^{\prime}x_{A}=B\varrho(p_{A})$
or
$\chi\geq\chi_{0}:=\underline{\chi}(t,p_{A}).$
Proof.
Technology is adopted if and only if the technology point $y=B\chi t$ lies on the profitable side of the supporting hyperplane through the autarky allocation $x_{A}$ with normal vector $p_{A}$. This is equivalent to $p_{A}^{\prime}y>p_{A}^{\prime}x_{A}$. ∎
The adoption condition does not depend on the overall scale of the budget $B$: both the value of the technology point ($p_{A}^{\prime}\,B\chi t$) and the value of the autarky allocation ($p_{A}^{\prime}x_{A}=B\varrho(p_{A})$) scale linearly in $B$, so $B$ cancels from the inequality. We interpret $\chi_{0}$ as the capability threshold for adopting some technology: at $\chi=\chi_{0}$ the worker is indifferent, and for $\chi$ just above $\chi_{0}$ optimal adoption can be arbitrarily small (the worker is at the extensive margin with $\lambda^{*}>0$ but close to zero).
Since $\cos\phi=p_{A}^{\prime}t$, the adoption condition becomes $\chi\cos\phi\geq\varrho(p_{A})$. The condition has a team-production interpretation: the collaboration is worthwhile when the AI’s contribution per unit of supervision ($\chi$), weighted by alignment ($\cos\phi$), exceeds the opportunity cost of supervision ($\varrho(p_{A})$). This implies a geometric representation of the adoption decision. We can invert the problem and ask: for a given technology capability $\chi$, what directions would the worker adopt?
Theorem 2.6 (Cone of technology adoption).
For a worker with autarky prices $p_{A}$ and unit revenue function $\varrho$, the cone of technology adoption for capability $\chi>\varrho(p_{A})$ is the set of technology directions the worker adopts:
$\mathcal{C}(\chi,p_{A},\varrho)=\left\{t\in\mathbb{R}^{N}:p_{A}^{\prime}t>\varrho(p_{A})/\chi,\quad\|t\|_{2}=1\right\}.$
The cone is centered on $p_{A}$ with half-angle $\phi_{0}=\arccos[\varrho(p_{A})/\chi]$. A worker adopts technology $(\chi,t)$ if and only if $t$ is within the cone of technology adoption, $t\in\mathcal{C}(\chi,p_{A},\varrho)$. Equivalently, the alignment angle $\phi$ is less than $\phi_{0}$.
The cone has several important properties. First, it is non-empty exactly when $\chi>\varrho(p_{A})$. When $\chi=\varrho(p_{A})$, only the direction $t=p_{A}$ is adopted. This provides a threshold condition for technology adoption that is distinct from absolute advantage. Near the threshold, the cone widens rapidly. Formally, $\partial\phi_{0}/\partial\chi$ is infinite at $\chi=\varrho$. Finally, as technology capability increases without bound, $\chi\to\infty$, the cone expands to a hemisphere ($\phi_{0}\to\pi/2$). Figure 5 illustrates the cone geometry and Figure 6 shows how the half-angle varies with capability.
$x_{1}$$x_{2}$$x_{3}$$x_{t}$ttOO
Figure 5: Cone of technology adoption in 3D
The unit sphere represents all possible technology directions $t$. The cone is centered on the autarky price $p_{A}$ with half-angle $\phi_{0}=\arccos(\varrho/\chi)$. The worker adopts any technology direction inside the cone. The dashed arcs show cross-sections of the cone at different capabilities $\chi$, the cross-section has radius $\sqrt{\chi^{2}-\varrho^{2}}$.
[…]
Figure 6: Cone half-angle as a function of technology capability
Corollary 2.1 (Comparative statics of the cone).
Let $\varrho=\varrho(p_{A})$ denote the worker’s unit revenue at autarky. The cone of technology adoption has the following properties as a function of capability $\chi$:
  1. 1.
    Threshold above absolute advantage. The cone is non-empty if and only if $\chi>\varrho$. Since $\varrho\geq 1$ generically, with equality only when the worker’s PPF passes through the unit simplex, adoption requires capability strictly above the absolute-advantage threshold $\chi=1$.
  2. 2.
    Fast widening near threshold. As $\chi\to\varrho^{+}$, the cone half-angle satisfies
    $\phi_{0}(\chi)\approx\sqrt{2\left(\frac{\chi}{\varrho}-1\right)},$
    so that $\phi_{0}$ grows with the square root of excess capability. In particular, $d\phi_{0}/d\chi\to+\infty$ as $\chi\to\varrho^{+}$: the range of adopted directions expands rapidly just above the threshold.
  3. 3.
    Monotonic widening. For all $\chi>\varrho$, $d\phi_{0}/d\chi>0$. The cone half-angle increases monotonically from 0 (at $\chi=\varrho$) to $\pi/2$ (as $\chi\to\infty$).
Figure 6 illustrates these properties: the half-angle is zero at $\chi=\varrho$, rises steeply just above the threshold (reflecting the square-root behavior), and increases monotonically toward $\pi/2$ as capability grows. The steep initial slope means that small improvements in capability near the threshold translate into large expansions of the set of adopted directions—a prediction that could help explain rapid adoption cascades when a technology crosses from “not quite good enough” to “just good enough.”
2.3  The intensive margin of technology
The cone of technology adoption only describes that the worker uses some technology, not how much. To understand this, we must also look at the production function $F$.
The intensive margin is economically central because it determines whether technology is a marginal supplement to human work or a full replacement. In particular, whether adoption eventually reaches $\lambda^{*}=1$ governs how much of the worker’s output is effectively pinned down by the technology direction rather than the worker’s own comparative advantage, and it determines how strongly technology reshapes the job’s implied shadow prices.
Let $y=\chi Bt$ the vector of activities that technology can produce when all the resources are spent on it. Optimal production is the solution to the problem
$F(X_{A},y):=\max_{x,\lambda}F((1-\lambda)x+\lambda y)\qquad\text{s.t.}\qquad x\in X_{A}.$
Under what conditions will $\lambda$ be interior, $\lambda\in(0,1)$?
Proposition 2.7 (Unique optimal technology use).
Suppose $F$ is strictly concave and $C^{2}$. Let $y=B\chi t$ and $X_{T}=\mathrm{conv}(X_{A}\cup\{y\})$. Then:
  1. 1.
[…]
  2. 2.
    There exist unique $x_{*}\in X_{A}$ and $\lambda_{*}\in[0,1]$ such that $\tilde{x}_{*}=(1-\lambda_{*})x_{*}+\lambda_{*}y$.
  3. 3.
[…]
Proof.
Uniqueness of $\tilde{x}_{*}$ follows from strict concavity of $F$ over the convex set $X_{T}$. The decomposition follows from the convex hull representation (Proposition 2.3). Uniqueness of $\lambda_{*}$ follows from strict concavity of $f(\lambda)=\max_{x\in X_{A}}F((1-\lambda)x+\lambda y)$, and $x_{*}$ is then pinned down by $x_{*}=(\tilde{x}_{*}-\lambda_{*}y)/(1-\lambda_{*})$ when $\lambda_{*}<1$. Uniqueness of $p_{100}$ follows from $C^{2}$ differentiability of $F$. See Appendix A.1. ∎
Take the optimal human task vector $x_{*}$. Connect this point with the point $y$ and study how the production function $F$ varies along this segment. Define
$f(\lambda)=F[(1-\lambda)x_{*}+\lambda y]\qquad\forall\lambda\in[0,1].$
This univariate function is strictly concave in $\lambda$ and maximized at $\lambda_{*}$. The solution is interior if and only if the derivative of $f$ is positive at zero and negative at one, $f^{\prime}(0)>0$ and $f^{\prime}(1)<0$.
The entry and corner conditions can be expressed in terms of the adoption threshold (Definition 2.8). See Appendix A.2 for the full derivation. The entry threshold and the corner threshold are both instances of the adoption threshold (Definition 2.8), evaluated at different prices:
$\chi_{0}=\underline{\chi}(t,p_{A}),\qquad\chi_{100}=\underline{\chi}(t,p_{100}),$
where $p_{100}=\nabla F(t)/\|\nabla F(t)\|$ is the shadow price at the technology direction.
Theorem 2.8 (Interior adoption).
Suppose $F$ is strictly concave, $C^{2}$, and homogeneous of degree one. Let $\lambda_{*}(\chi)$ denote the unique optimal technology intensity from Proposition 2.7. Then:
  1. 1.
    No adoption: If $p_{A}^{\prime}t=0$, then $\lambda_{*}=0$ for all $\chi$: the worker never adopts.
  2. 2.
    Never all-in: If $F(t)=0$, then $\lambda_{*}<1$ for all $\chi$: the worker never goes all in.
  3. 3.
    Partial adoption: If $p_{A}^{\prime}t>0$ and $F(t)>0$, then $\chi_{100}>\chi_{0}$: partial adoption occurs for $\chi\in(\chi_{0},\chi_{100})$, with $\lambda_{*}\in(0,1)$. For $\chi\geq\chi_{100}$, $\lambda_{*}=1$ (all-in).
Proof sketch.
Cases 1–2 follow from the boundary conditions of the univariate function $f(\lambda)=F[(1-\lambda)x_{*}+\lambda y]$. For case 3, the strict inequality $\chi_{100}>\chi_{0}$ follows from strict concavity: the supporting hyperplanes at $x_{A}$ and at the technology direction $t$ are distinct whenever $t$ is not collinear with $x_{A}$. See Appendix A.2. ∎
To compute these thresholds in practice, one proceeds in two steps. First, solve the autarky problem to obtain the autarky allocation $x_{A}$ and its supporting normal $p_{A}$ (under differentiability, $p_{A}\propto\nabla g(x_{A})$, normalized to unit length). Given $p_{A}$, compute $\chi_{0}=\underline{\chi}(t,p_{A})=\varrho(p_{A})/(p_{A}^{\prime}t)$. Second, compute the technology-induced price $p_{100}$ from the job’s marginal rates of substitution at the technology direction (under differentiability, $p_{100}\propto\nabla F(t)$, normalized). Then compute $\chi_{100}=\underline{\chi}(t,p_{100})$. The gap $\chi_{100}-\chi_{0}>0$ determines the width of the partial-adoption region.
Figure 7 illustrates how to compute $\chi_{0}$ and $\chi_{100}$ in two dimensions. At capability $\chi_{0}$ the technology vector is just sufficiently long to touch the budget line (hyperplane in $N$ dimensions) spanned by the autarky prices at the production point $x_{A}$. As capability grows, it will eventually reach a point where the production isoquant is parallel to the PPF at a new point $x_{100}$. At this capability, production is maximized by using technology 100 percent of the time of the worker. As capability increases, this corner solution is going to remain, with the slope of the isoquant no longer pinning down technology usage.
The gap between $\chi_{100}$ and $\chi_{0}$ determines how long is the range of the “intensive margin” of technology usage. This depends on the curvature of both $g$ and $F$. For general curved PPFs, $x_{100}$ is typically distinct from $x_{A}$ and is “directionally farther” from the technology. This means that the last few human tasks that are done before full AI adoption are very different from what AI has to offer.
[…]
Figure 7: The limits of interior technology adoption
For technology adoption, the worker’s skills and job requirements matter only through shadow prices. The autarky prices $p_{A}$ summarize which task bundles are scarce at the margin in the current worker–job match: they encode the relevant interaction between the resource use function $g$ (skills) and the production function $F$ (requirements). The adoption condition in Theorem 2.5 depends on $g$ and $F$ only through $p_{A}$ and the induced unit revenue $\varrho(p_{A})$, and the all-in threshold depends on $F$ only through $p_{100}$. As a result, two matches that generate the same $p_{A}$ face the same cone of adopted technology directions for any given capability $\chi$.
This sufficient-statistic property is the reason the worker–job match is the relevant unit of analysis and why one-dimensional exposure or skill rankings can misclassify adoption: what matters is the direction of shadow prices, not a scalar summary. We return to this interpretation in the Introduction.
2.4  More than one technology
Workers rarely face a single tool in isolation: multiple technologies coexist, and adopting one reshapes the feasible set and the marginal value of tasks. This matters because adoption decisions are value comparisons at shadow prices. When a worker adopts an initial technology, the implied prices rotate, which can make a second technology more or less attractive even if its direction and capability are unchanged. This section extends the adoption and shadow-price logic to settings where several directional technologies are available simultaneously.
We proceed by induction on the number of technologies. The key insight is that after adopting $K$ technologies, the expanded feasible set $X_{T,K}$ retains the properties of the autarky PPS—convexity, compactness, free disposal—so the adoption logic for technology $K+1$ is identical in structure to the single-technology case, with $X_{T,K}$ replacing $X_{A}$ and the current shadow prices $p_{K}^{*}$ replacing $p_{A}$.
Definition 2.9 (Multi-technology PPS).
Given technologies $(t_{1},\chi_{1}),\ldots,(t_{K},\chi_{K})$ with technology points $y_{k}=B\chi_{k}t_{k}$, the $K$-technology PPS is
$X_{T,K}=\left\{z\in\mathbb{R}^{N}_{+}:z\leq x_{H}+\sum_{k=1}^{K}\lambda_{k}y_{k},\;g(x_{H})\leq\left(1-\sum_{k=1}^{K}\lambda_{k}\right)B,\;\lambda_{k}\geq 0,\;\sum_{k=1}^{K}\lambda_{k}\leq 1\right\}.$
Under our assumptions on $g$, $X_{T,K}=\text{conv}(X_{A}\cup\{y_{1},\ldots,y_{K}\})$ with free disposal (Proposition 2.3 applied inductively). See Appendix A.3.
Proposition 2.9 (Inductive adoption structure).
Let $\tilde{x}_{K}^{*}$ denote the optimum over $X_{T,K}$ and $p_{K}^{*}=\nabla F(\tilde{x}_{K}^{*})/\|\nabla F(\tilde{x}_{K}^{*})\|$ the associated shadow price. Then:
  1. 1.
    Entry: Technology $K+1$ with direction $t_{K+1}$ is adopted if and only if $\chi_{K+1}>\underline{\chi}(t_{K+1},p_{K}^{*})$, where $\underline{\chi}(t,p)=\varrho_{K}(p)/(p^{\prime}t)$ and $\varrho_{K}(p)=\max_{z\in X_{T,K}}p^{\prime}z/B$ is the unit revenue over the $K$-technology PPS.
  2. 2.
    Rising bar: $\varrho_{K}(p)\geq\varrho_{K-1}(p)$ for all $p$ (the maximum over a larger set is weakly larger), with strict inequality for any $p$ such that one of $y_{1},\ldots,y_{K}$ lies on the profitable side of the supporting hyperplane at $p$. Hence the entry threshold for technology $K+1$ is weakly higher than for technology $K$: each successive technology faces a tougher adoption bar.
  3. 3.
    All-in: The worker goes all-in on technology $K+1$ if $\chi_{K+1}>\varrho_{K}(p_{100,K+1})/p_{100,K+1}^{\prime}t_{K+1}$ , and $\chi_{K+1}/\chi_{k}>p_{100,K+1}^{\prime}t_{k}^{\prime}p_{100,K+1}^{\prime}t_{1}$ for all k=1,2,,Kk=1,2,...,K, where $p_{100,K+1}=\nabla F(t_{K+1})/\|\nabla F(t_{K+1})\|$.
Proof.
Part 1 follows from applying Theorem 2.5 with $X_{T,K}$ in place of $X_{A}$: the entry condition is $p_{K}^{*\prime}y_{K+1}>p_{K}^{*\prime}\tilde{x}_{K}^{*}=B\varrho_{K}(p_{K}^{*})$. Part 2: $X_{T,K}\supseteq X_{T,K-1}$, so the support function (maximum over feasible bundles) is weakly larger; strictly so when the new technology point expands the set in direction $p$. Part 3 applies the all-in dominance logic of Theorem 2.8 to $X_{T,K}$: at prices $p_{100,K+1}$, no element of $X_{T,K}$—neither human bundles nor previous technology points—delivers higher value. See Appendix A.4. ∎
The all-in condition has a simple interpretation. Evaluate all feasible alternatives at the marginal value of output in the neighborhood of going all-in on the new tool, summarized by $p_{100,K+1}$. The worker chooses $y_{K+1}$ if, at these prices, the new technology delivers strictly higher value than any purely human bundle and than any previously available technology point. In other words, near $y_{K+1}$ there is no competitive human or older-technology substitute.
3  A parametric illustration
This section provides a tractable parametric specialization of the general model. We model job requirements with a constant elasticity of substitution (CES) production function and skills with a constant elasticity of transformation (CET) resource use function. CES means that tasks are substitutable inputs in producing output, with elasticity parameter $\sigma$. CET means that the worker can transform a fixed resource budget into different task bundles with constant elasticity $\gamma$; it is a convenient parametric way to represent the Production Possibility Set.
The goal of the CES–CET specialization is to make the autarky shadow-price vector $p^{A}$ operational by expressing it as a function of primitives: job requirements $\theta$ and skills $s$. Because adoption decisions depend on $p^{A}$ (Theorems 2.5 and 2.6), this mapping turns the qualitative geometry of Section 2 into closed-form comparative statics.
3.1  CES job requirements and CET skills
Definition 3.1 (CES–CET parametric environment).
The CES–CET environment specializes the general model (Assumptions 2.1 and 2.2) as follows.
Job requirements (CES). The production function is constant elasticity of substitution (CES):
$F(x)=\left(\sum_{i=1}^{N}\theta_{i}^{1/\sigma}\,x_{i}^{(\sigma-1)/\sigma}\right)^{\sigma/(\sigma-1)},$
where $\theta=(\theta_{1},\ldots,\theta_{N})$ with $\theta_{i}>0$ captures job requirements (task $i$’s value weight) and $\sigma>0$ is the elasticity of substitution across tasks. $F$ satisfies constant returns, strict quasi-concavity, and $C^{2}$ differentiability on $\mathbb{R}^{N}_{++}$.
Skills (CET). The resource use function is constant elasticity of transformation (CET):
$g(x)=\left(\sum_{i=1}^{N}s_{i}^{-1/\gamma}\,x_{i}^{(\gamma+1)/\gamma}\right)^{\gamma/(\gamma+1)},$
where $s=(s_{1},\ldots,s_{N})$ with $s_{i}>0$ captures skills (higher $s_{i}$ means lower resource cost for task $i$) and $\gamma>0$ is the elasticity of transformation across tasks. $g$ is convex, homogeneous of degree one, increasing, and $g(0)=0$. See Appendix A.5.
The parameters have straightforward economic content. The vector $\theta$ captures job requirements: higher $\theta_{i}$ means task $i$ is more valuable for producing output. The vector $s$ captures skills: higher $s_{i}$ means task $i$ uses fewer resources at the margin. The curvature parameters capture flexibility. A higher substitution elasticity $\sigma$ makes tasks more interchangeable in production (a more flexible job), while a higher transformation elasticity $\gamma$ makes it easier for the worker to reallocate effort across tasks (a more flexible skill set).
3.2  Autarky allocation in closed form
We begin with the autarky allocation problem (Definition 2.1). Under the CES–CET environment, the optimizer admits a closed-form expression. Define the CES–CET productivity index
$\Phi(\theta,s)=\sum_{j=1}^{N}s_{j}^{(\sigma-1)/(\gamma+\sigma)}\,\theta_{j}^{(\gamma+1)/(\gamma+\sigma)}.$
Then the autarky allocation takes the form
$x_{i}^{A}=B\cdot\frac{s_{i}^{\gamma/(\gamma+\sigma)}\,\theta_{i}^{\sigma/(\gamma+\sigma)}}{\Phi^{\sigma/(\gamma+\sigma)}},\qquad i=1,\ldots,N,$
and autarky output per unit budget satisfies
$\frac{Y^{A}}{B}=\frac{F(x^{A})}{B}=\Phi^{\sigma/(\sigma-1)}.$
See Appendix A.5 for derivations.
The allocation formula shows that the worker concentrates activity where skills and requirements overlap. The exponents make the role of curvature transparent. When $\sigma$ is high, tasks are easier to substitute in production, so the worker tilts more strongly toward high-requirement activities (higher weight on $\theta_{i}$). When $\gamma$ is high, it is easier to reallocate effort across tasks inside the constraint set, so differences in skills translate more strongly into differences in chosen activity levels.
The productivity index $\Phi$ has a simple alignment interpretation: output is high when the worker is relatively skilled in the tasks that the job values (roughly, when $s_{i}$ is high where $\theta_{i}$ is high). Misalignment lowers $\Phi$ by forcing the worker to spend scarce resources on tasks that are both required and expensive to execute.
3.3  Shadow prices as functions of job requirements and skills
In the general model, autarky prices $p^{A}$ are defined by tangency between the Production Possibility Frontier and an isoquant. Under CES–CET, they take a simple power-law form:
$p_{i}^{A}\propto\left(\frac{\theta_{i}}{s_{i}}\right)^{1/(\gamma+\sigma)},\qquad\|p^{A}\|_{2}=1.$
See Appendix A.5.
This formula clarifies the economic content of shadow prices. A task has high marginal value when it is important for output (high $\theta_{i}$) and costly to execute for the worker (low $s_{i}$). The relevant object is therefore the ratio $\theta_{i}/s_{i}$: the gap between requirements and ease of execution.
The distinction between $x^{A}$ and $p^{A}$ is central for interpreting observed behavior. A worker may spend substantial time on a task either because it is required (high $\theta_{i}$) or because it is easy (high $s_{i}$). Shadow prices separate these forces: $p_{i}^{A}$ is high precisely when the job demands the task but the worker finds it expensive. This is the key statistic for valuing a directional technology.
3.4  Comparative statics inside the job (Jevons Paradox)
The closed forms make it easy to describe within-job reallocation. Define the autarky time (or activity) share of task $i$ as
$\omega_{i}=\frac{s_{i}^{-1/\gamma}x_{i}^{1+1/\gamma}}{\sum_{j=1}^{N}s_{j}^{-1/\gamma}x_{j}^{1+1/\gamma}}.$
Substituting the allocation formula above, all common scaling terms cancel and we obtain
$\omega_{i}^{A}=\frac{s_{i}^{(\sigma-1)/(\gamma+\sigma)}\,\theta_{i}^{(\gamma+1)/(\gamma+\sigma)}}{\Phi}.$
An increase in skill $s_{i}$ raises $\omega_{i}^{A}$ if and only if the tasks are substitutes in the production function, $\sigma>1$. This is Jevons’ Paradox for knowledge work. Someone skilled at a task will spend more time with it, not less, when demand for that task is elastic. By contrast, when tasks are complement, each task has to be completed to some predefined level, and skilled workers spend less time on satisfying the job requirement. This also highlights that activities, in general, are poor measures of skills. ## Applying the adoption theorems under CES–CET
We now express the key statements about technology adoption in this parametric environment. Theorems 2.5 and 2.6 show that adoption depends on (i) autarky prices $p^{A}$ (through alignment $p^{A}\cdot t$) and (ii) the unit revenue function $\varrho(p^{A})$. Under CES–CET, both objects can be expressed in closed form as functions of $(\theta,s,\sigma,\gamma)$.
Proposition 3.1 (Adoption objects under CES–CET).
Under the CES–CET environment (Definition 3.1), let $\Phi=\sum_{j}s_{j}^{(\sigma-1)/(\gamma+\sigma)}\theta_{j}^{(\gamma+1)/(\gamma+\sigma)}$:
  1. 1.
    Unit revenue. The unit revenue function is the dual of the CET:
    $\varrho(p)=\left(\sum_{i=1}^{N}s_{i}\,p_{i}^{\gamma+1}\right)^{1/(\gamma+1)}.$
  2. 2.
    Entry threshold. In terms of primitives $(\theta,s,\sigma,\gamma,t)$:
    $\chi_{0}(t)=\frac{\Phi^{1/(\gamma+1)}}{\sum_{i}t_{i}(\theta_{i}/s_{i})^{1/(\gamma+\sigma)}}.$
    The numerator $\Phi^{1/(\gamma+1)}$ measures the worker’s autarky productivity: it is high when skills align with requirements. The denominator measures alignment of autarky prices with the technology direction: it is high when $t_{i}$ loads on tasks where $\theta_{i}/s_{i}$ is large (required but hard). Better-aligned technology faces a lower entry bar.
  3. 3.
    Corner threshold. The corner threshold can be written as
    $\chi_{100}(t)=\frac{\left(\sum_{i}s_{i}\,\theta_{i}^{(\gamma+1)/\sigma}\,t_{i}^{-(\gamma+1)/\sigma}\right)^{1/(\gamma+1)}}{\sum_{i}\theta_{i}^{1/\sigma}\,t_{i}^{1-1/\sigma}}.$
    The numerator depends on $(s,\theta,\sigma,\gamma,t)$: the worker’s unit revenue at technology-induced prices. The denominator depends on $(\theta,\sigma,t)$ only: how much value the technology direction generates at its own induced prices. Skills $s$ enter only the numerator: a more skilled worker has higher $\chi_{100}$ (harder to push to all-in).
See Appendix A.5 for derivations.
The next corollary derives some comparative statics of the adoption thresholds with respect to the curvature of the PPF and the production function. The first statement says that when jobs or skills are more flexible, even a less aligned technology can be productive. Adoption is faster in flexible jobs like general knowledge work (e.g., software, consulting, academia, journalism), and slower in very rigid jobs requiring highly specialized skills (e.g. medicine) or exact satisfaction of job requirements (e.g., law, medicine, nuclear physics, aviation). The second statement says that flexible jobs not only adopt early, but are also more likely to go “all in.” For more rigid jobs, we will see a wider range of interior AI usage with $\lambda\in(0,1)$.
Corollary 3.1 (Curvature and adoption).
Fix autarky productivity $\Phi$, job requirements $\theta$, skills $s$, and technology capability $\chi$. Then:
  1. 1.
    Cone width increases with flexibility. The cone of adopted technology directions is wider when $\gamma+\sigma$ is large. In the limit $\gamma+\sigma\to\infty$, the cone converges to a hemisphere (all directions with positive alignment are adopted).
  2. 2.
    Partial-adoption region narrows with flexibility. The ratio $\chi_{100}/\chi_{0}$ is decreasing in $\gamma$ and in $\sigma$. In the limit, adoption becomes all-or-nothing: either $\lambda_{*}=0$ or $\lambda_{*}=1$.
Proof sketch.
(1) The autarky price satisfies $p_{i}^{A}\propto(\theta_{i}/s_{i})^{1/(\gamma+\sigma)}$. As $\gamma+\sigma\to\infty$, the exponent $1/(\gamma+\sigma)\to 0$, so $p^{A}$ converges to the uniform direction. A more uniform price vector is positively aligned with a larger set of technology directions, widening the cone. (2) The thresholds $\chi_{0}$ and $\chi_{100}$ are evaluated at $p_{A}$ and $p_{100}$ respectively. Higher curvature compresses both price vectors toward uniformity, reducing the gap between them. See Appendix A.6. ∎
4  Discussion and Relation to the Literature
Task-based labor-market models have a long history, and our framework builds on several strands of this literature. To clarify what is new, we organize prior work along five dimensions: (i) job heterogeneity (which task directions matter across jobs); (ii) worker task elasticity $\gamma$ (how flexibly workers reallocate across tasks); (iii) production elasticity $\sigma$ (how tasks combine in output); (iv) skill heterogeneity (how workers differ in task-specific productivity); and (v) technology directionality (whether technical change favors particular task directions).
4.1  Organizing Framework
Our model has four primitives: job requirements $\theta\in\mathbb{R}^{N}_{+}$, worker skills $s\in\mathbb{R}^{N}_{+}$, transformation elasticity $\gamma>0$ (the curvature of the worker’s task-allocation frontier), and substitution elasticity $\sigma>0$ (the curvature of the production aggregator). Technology enters as a direction $t\in S^{N-1}$ and capability $\chi>0$. This nests several important special cases.
Table 1: Task-based models: key parameters
Paper Jobs $\gamma$ $\sigma$ Workers Technology
Welch (1969) NN-dim 0 $F$ NN-dim
Katz and Murphy (1992) NN-dim 0 $F$ NN-dim general
—application 2 types 0 $\sigma$ 2 types factor-augmenting
Autor et al. (2003) 2 tasks $\infty$ 1 1-dim routine-biased
Lazear (2009) 2 skills 0 $\infty$ 2-dim
Acemoglu and Autor (2011) continuum $\infty$ 1 3 types task threshold
Firpo et al. (2011) 5 tasks $\infty$ >1>1 5-dim 5-dim
Cortes (2016) 3 tasks $\infty$ 0 routine-biased
Fortin and Lemieux (2016) NN-dim 0 $\infty$ NN-dim
Acemoglu and Restrepo (2018) continuum $\infty$ $\sigma$ 1 type rotating threshold
Acemoglu and Restrepo (2020) continuum $\infty$ 1 1 type task threshold
Bárány et al. (2020) NN-dim 0 $\infty$ NN-dim
Böhm (2020) $K$ tasks logit $\infty$ routine-biased
—application 3 tasks logit $\infty$ routine-biased
Cavounidis and Lang (2020) NN-dim $\sigma$ $\infty$ NN-dim
Cavounidis et al. (2024) NN-dim $\infty$ $\sigma$ NN-dim within-occ
Autor and Thompson (2025) 1-dim $\infty$ 0 1-dim 1-dim
Freund and Mann (2025) NN-dim $\infty$ 1 NN-dim directional
Althoff and Reichardt (2026) NN-dim $\infty$ $\sigma$ directional
Benzell and Myers (2026) NN-dim 0 $\sigma$ NN-dim 1-dim
Gans and Goldfarb (2026) $n$ tasks $\infty$ 1 1-dim
Bárány and Holzheu (2026) NN-dim 0 quadratic NN-dim
This paper NN-dim $\gamma$ $\sigma$ NN-dim directional
Note: “Jobs” and “Workers” refer to the dimensionality of heterogeneity in job requirements and worker skills, respectively. This is distinct from the number of tasks $N$: a model with many tasks can still be one-dimensional if tasks are ordered by a single attribute (e.g., routine content, expertise level).
Many models take corner cases on at least one of the two curvature margins. It is common to assume either Leontief aggregation ($\sigma=0$) or perfect substitution ($\sigma=\infty$), and either rigid task endowments ($\gamma=0$) or frictionless reallocation across tasks ($\gamma=\infty$). Allowing both $\gamma$ and $\sigma$ to be finite matters because shadow prices and the direction of marginal value depend on both worker flexibility and production complementarity. This is the key step behind our alignment-based adoption condition and the structured intensive margin.
4.2  Multidimensional Heterogeneity
A key feature is that jobs and workers differ along multiple dimensions that cannot be collapsed to a single index. This rules out simple rankings of “high-skill” versus “low-skill” or “routine” versus “nonroutine” as sufficient statistics.
Welch (1969) is among the first to decompose labor into multiple “labor inputs”—what we now call tasks. Workers have fixed endowments of these inputs ($\gamma=0$), and a general production function $F$ aggregates them. This early framework already allows for $N$-dimensional heterogeneity in both jobs and workers.
Firpo et al. (2011) take a step toward higher dimensionality, using five O*NET task measures. Their reduced-form specification writes wages as an inner product of skills and returns—analogous to our $\langle s,\theta\rangle$ structure. Importantly, because task prices vary across occupations, tasks are implicitly not fully traded ($\gamma<\infty$). Their Oaxaca-Blinder decomposition captures changes in task content across occupations, which corresponds to “rotation” in our framework.
Cortes (2016) studies job polarization with three task types and Leontief production ($\sigma=0$). Workers choose tasks freely ($\gamma=\infty$), and routine-biased technical change shifts employment across task categories. Böhm (2020) estimates task prices using a generalized Roy model where workers sort across $K$ tasks via discrete choice (effectively $\gamma=\text{logit}$). Tasks are traded at economy-wide prices ($\sigma=\infty$), and the focus is on recovering the price effects of routine-biased technical change.
Two recent papers treat skills as genuinely multidimensional objects rather than as a scalar index. Bárány et al. (2020) develop and estimate a multi-sector general-equilibrium model of education and sector-occupation choices; jobs are bundles of tasks and workers sort on a multidimensional skill vector that evolves with schooling and experience (Abstract). Bárány and Holzheu (2026) define worker specialization as the distance between a worker’s skill set and those prevalent in the labor market and study its wage-unemployment trade-off in a random-search model with mismatch penalties and skill complementarity (Abstract). Relative to our framework, these papers emphasize schooling, sorting, and search frictions, whereas we focus on within-job task choice and the alignment-based adoption decision.
Benzell and Myers (2026) is among the first to model within-job production with high-dimensional, non-traded tasks. Workers have skill endowments ($\gamma=0$), and tasks combine with CES elasticity $\sigma$. Automation supersedes one task direction. They emphasize that in high dimensions, workers cannot be ranked by a single index. Our model generalizes by allowing $\gamma>0$ (workers can reallocate across tasks) and directional technology.
4.3  Within-Job Task Reallocation
A second feature of our model is within-job task reallocation: workers shift effort across tasks while holding the job fixed. This margin is central to understanding how AI changes work without changing employment.
Lazear (2009) introduces skills with job-specific weights—a direct antecedent of our $(\theta,s)$ formulation. In his baseline model, workers have endowments of two skills ($\gamma=0$), and firms combine them linearly with heterogeneous weights ($\sigma=\infty$). All skills are general, but firms value them differently, generating apparent firm-specific human capital. The model does not include technology, but the skill-weights structure is foundational. Fortin and Lemieux (2016) extend this skill-weights approach to study inequality, combining it with the task framework of Acemoglu and Autor (2011).
Spitz-Oener (2006) provides early measurement evidence that task content changes within occupations over time, with computer use associated with shifts toward nonroutine cognitive tasks. Cavounidis and Lang (2020) microfound dynamic skill formation in a multidimensional setting. In their static model, workers choose task bundles with elasticity $\gamma=\sigma$ (in our notation) and output sums tasks linearly ($\sigma=\infty$), with job-specific task weights corresponding to our $\theta$. Cavounidis et al. (2024) apply this framework empirically, documenting within-occupation shifts in task content (what we call “rotation”) and connecting productivity effects to an envelope condition (our shadow prices).
Autor and Thompson (2025) model expertise as a scalar that determines how task bundles translate into output. Technology compresses the expertise distribution by making low-expertise workers more productive. The task space is implicitly high-dimensional but collapsed to a single index. Our framework preserves the full dimensionality, showing that scalar indices can mask important heterogeneity in alignment.
Freund and Mann (2025) develop a model where workers have multidimensional skill portfolios and jobs bundle multiple tasks, using 38 O*NET task clusters in their application. Tasks are freely tradable ($\gamma=\infty$) and combine with Cobb-Douglas technology ($\sigma=1$). They emphasize alignment between worker skills and job requirements and show that one-dimensional “exposure” measures can miss important heterogeneity. Their wage formula resembles our productivity expression, and their notion of skill-job “correlation” parallels our inner product structure. Our framework generalizes by allowing finite $\gamma$ (non-traded tasks) and finite $\sigma$ (task complementarity).
Althoff and Reichardt (2026) develop a closely related framework where workers allocate time across $N$ tasks ($\gamma=\infty$) and output is CES with elasticity $\rho$. Technical change is task-specific and directional, allowing them to study how AI shifts comparative advantage across occupations. They extend the analysis to occupational choice, showing how task-specific technology shocks induce reallocation across jobs.
Gans and Goldfarb (2026) study an O-ring production function where workers allocate time across $N$ tasks. Automation makes workers redundant in specific tasks, but reallocation occurs within the job. This is close to our setup with $\gamma=\infty$ and $\sigma=1$. The key difference is that their technology is one-dimensional (automation of an adjacent range of tasks), whereas ours is directional (technology produces a bundle of tasks in fixed proportions).
Lin (2025) provides direct evidence for within-job task reallocation in response to AI changes. Using prompt-level data from Midjourney, she shows that most of the output shift after an upgrade comes from users changing their prompts rather than from the capability change itself. This is evidence of rotation in response to changing AI capabilities, exactly the intensive margin our model emphasizes.
4.4  Other Task-Based Models
Several other papers develop task-based frameworks that inform our approach but emphasize different margins.
Katz and Murphy (1992) extend the Welch (1969) framework to study wage inequality: their general model allows $N$ labor inputs with a generic concave production function; their empirical application collapses to two skill types (college/non-college) with CES aggregation ($\sigma\approx 1.4$). They treat technical change as shifts in labor demand without specifying direction.
The skill-biased technical change literature (Katz and Murphy, 1992; Acemoglu, 2002) partitions workers into broad skill groups and models technology as factor-augmenting. This approach captures between-group inequality but cannot address within-occupation heterogeneity in adoption or gains. The routine-biased framework (Autor et al., 2003; Acemoglu and Autor, 2011) adds a task dimension: technology substitutes for routine tasks while complementing non-routine tasks. Workers choose task bundles, but the task space is one-dimensional (routine vs. non-routine), and tasks are traded at economy-wide prices ($\gamma=\infty$).
Acemoglu and Restrepo (2018, 2020) develop a workhorse model where a continuum of tasks is traded at economy-wide prices. Automation shifts the threshold between tasks performed by labor and capital. The production elasticity $\sigma$ governs macroeconomic implications. Our framework nests this as the limiting case $\gamma\to\infty$ (tasks fully tradable) with technology “rotating” the threshold across task indices.
Lise and Postel-Vinay (2020) develop a sorting model with multidimensional skills. Workers and jobs match in a frictional labor market. The focus is on sorting and human capital accumulation rather than technology adoption.
Agrawal et al. (2026) emphasize that AI may relax bottlenecks rather than automate tasks. This is complementary to our approach: in our model, AI that produces tasks where the worker is constrained (high shadow price) relaxes the binding constraint, which is the geometric interpretation of bottleneck relief.
The contribution of our framework is to combine five features that prior models typically treat separately: (i) a high-dimensional task space; (ii) finite worker task elasticity $\gamma$ (non-traded tasks and worker-specific shadow prices); (iii) finite production elasticity $\sigma$ (task complementarity); (iv) multidimensional skill heterogeneity $s$; and (v) directional technology. The interaction of these features yields the adoption cone, the alignment condition, and the intensive margin.
5  Conclusion
GenAI is an unusually easy-to-adopt general-purpose technology, but its usefulness varies sharply across directions in a high-dimensional task space. This directional heterogeneity makes one-dimensional notions of exposure or skill unreliable guides to either adoption or productivity gains.
We model a job as a neoclassical task allocation problem: workers choose task bundles subject to a convex resource constraint, and the optimal allocation is characterized by a vector of shadow (autarky) prices that summarize which tasks are scarce at the margin in the current worker-job match. We model AI as a directional technology with a capability level. Adoption then reduces to a value comparison at shadow prices.
Three implications follow. First, absolute advantage is necessary but not sufficient: even a superior tool may not be adopted when its strengths point away from what is valuable at the margin. Second, adoption has a sharp entry threshold; near it, small capability gains can produce large expansions in the set of adopted directions. Third, adoption has a structured intensive margin: because entry and all-in thresholds generally differ, optimal use can remain partial over a range of capabilities, generating stable hybrid production.
These mechanisms are match-specific because shadow prices depend jointly on worker skills and job requirements, and they extend naturally to settings with multiple tools, where adoption rotates shadow prices and can raise the bar for subsequent technologies.
Appendix A Proofs and Derivations
A.1  Proof of Proposition 2.7
[…]
Uniqueness of $\lambda_{*}$. Define $h(\lambda)=\max_{x\in X_{A}}F((1-\lambda)x+\lambda y)$. For each fixed $x\in X_{A}$, $\lambda\mapsto F((1-\lambda)x+\lambda y)$ is strictly concave (composition of strictly concave $F$ with an affine function of $\lambda$). The function $h(\lambda)$ is strictly concave by a standard parametric optimization argument: take $\lambda_{1}\neq\lambda_{2}$ with maximizers $x_{1}^{*},x_{2}^{*}$; the two optimal bundles $(1-\lambda_{i})x_{i}^{*}+\lambda_{i}y$ are distinct (since $\tilde{x}_{*}$ is unique and varies with $\lambda$), so strict concavity of $F$ gives $h(\bar{\lambda})>\mu h(\lambda_{1})+(1-\mu)h(\lambda_{2})$. Hence $\lambda_{*}$ is unique, and $x_{*}=(\tilde{x}_{*}-\lambda_{*}y)/(1-\lambda_{*})$ when $\lambda_{*}<1$.
[…]
A.2  Proof of Theorem 2.8
Define the univariate function $f(\lambda)=F[(1-\lambda)x_{*}+\lambda y]$, which is strictly concave in $\lambda$ on $[0,1]$.
Case 1. If $p_{A}^{\prime}t=0$, then $f^{\prime}(0)=\nabla F(x_{A})^{\prime}(y-x_{A})=B\chi\nabla F(x_{A})^{\prime}t-\nabla F(x_{A})^{\prime}x_{A}$. Since $\nabla F(x_{A})\propto p_{A}$ and $p_{A}^{\prime}t=0$, we have $\nabla F(x_{A})^{\prime}t=0$, so $f^{\prime}(0)<0$ and $\lambda_{*}=0$.
Case 2. If $F(t)=0$, then $F(y)=F(B\chi t)=B\chi F(t)=0$ by homogeneity. Going all-in yields zero output, while autarky gives positive output $F(x_{A})>0$. Hence $\lambda_{*}<1$.
Case 3. When $p_{A}^{\prime}t>0$ and $F(t)>0$, both $\chi_{0}=\underline{\chi}(t,p_{A})$ and $\chi_{100}=\underline{\chi}(t,p_{100})$ are finite. The entry condition $f^{\prime}(0)>0$ holds iff $\chi>\chi_{0}$, and the corner condition $f^{\prime}(1)<0$ holds iff $\chi<\chi_{100}$. By Lemma A.1 below, $\chi_{100}>\chi_{0}$ whenever $t$ is not collinear with $x_{A}$. For $\chi\in(\chi_{0},\chi_{100})$, we have $f^{\prime}(0)>0$ and $f^{\prime}(1)<0$, so by strict concavity $\lambda_{*}\in(0,1)$. For $\chi\geq\chi_{100}$, $f^{\prime}(1)\geq 0$, so $\lambda_{*}=1$. $\square$
Lemma A.1 (Threshold ordering).
If $F$ is strictly concave, homogeneous of degree one, and the technology direction $t$ is not collinear with the autarky allocation $x_{A}$, then $\chi_{100}>\chi_{0}$.
Proof.
Define $c=F(x_{A})/F(t)$, so that $ct$ lies on the same isoquant as $x_{A}$. By strict concavity, the supporting hyperplane at $x_{A}$ lies strictly above $F(t)$: $F(t)<\nabla F(x_{A})^{\prime}t$. Using Euler’s theorem $F(x_{A})=\nabla F(x_{A})^{\prime}x_{A}$ and the definition $\chi_{0}=F(x_{A})/(\nabla F(x_{A})^{\prime}t)$, we have $c>\chi_{0}$. Similarly, the supporting hyperplane at $t$ lies strictly above $F(x_{A})$: $F(x_{A})<\nabla F(t)^{\prime}x_{A}$. Since $\chi_{100}\geq(\nabla F(t)^{\prime}x_{A})/F(t)$ and $\nabla F(t)^{\prime}x_{A}>F(x_{A})$, we have $\chi_{100}>c$. Combining: $\chi_{0}<c<\chi_{100}$. $\square$
A.3  Multi-technology PPS
Proposition A.1.
Under our assumptions on $g$ (convex, homogeneous of degree 1), $X_{T,K}=\mathrm{conv}(X_{A}\cup\{y_{1},\ldots,y_{K}\})$.
Proof.
$X_{T,K}\subseteq\mathrm{conv}(X_{A}\cup\{y_{1},\ldots,y_{K}\})$:* Take $z=x_{H}+\sum_{k}\lambda_{k}y_{k}\in X_{T,K}$ with $\Lambda=\sum_{k}\lambda_{k}$. For $\Lambda<1$, define $\tilde{x}=x_{H}/(1-\Lambda)$. By homogeneity, $g(\tilde{x})=g(x_{H})/(1-\Lambda)\leq B$, so $\tilde{x}\in X_{A}$. Then $z=(1-\Lambda)\tilde{x}+\sum_{k}\lambda_{k}y_{k}$ is a convex combination.
$\mathrm{conv}(X_{A}\cup\{y_{1},\ldots,y_{K}\})\subseteq X_{T,K}$:* Take $z=\mu_{0}x+\sum_{k}\mu_{k}y_{k}$ with $x\in X_{A}$. Set $\lambda_{k}=\mu_{k}$ and $x_{H}=\mu_{0}x$. Then $g(x_{H})=\mu_{0}g(x)\leq\mu_{0}B=(1-\sum_{k}\lambda_{k})B$. $\square$
Lemma A.2 (Unique shadow prices).
If $F$ is strictly concave and $C^{1}$, the optimum $x^{*}_{K}=\arg\max_{x\in X_{T,K}}F(x)$ is unique and $p^{*}_{K}=\nabla F(x^{*}_{K})/\|\nabla F(x^{*}_{K})\|$ is well-defined.
Proof.
Uniqueness of $x^{*}_{K}$ follows from strict concavity of $F$ over the convex set $X_{T,K}$. Continuous differentiability of $F$ ensures $\nabla F(x^{*}_{K})$ exists and is nonzero, giving a unique $p^{*}_{K}$. $\square$
A.4  All-in adoption of technology $K+1$
Proposition A.2.
Under Assumption 2.2, the worker goes all-in on technology $K+1$ if
$p_{100,K+1}^{\prime}y_{K+1}>\max\left\{\max_{x\in X_{A}}p_{100,K+1}^{\prime}x,\;p_{100,K+1}^{\prime}y_{1},\ldots,p_{100,K+1}^{\prime}y_{K}\right\}.$
Proof.
At $x^{*}=y_{K+1}$, the first-order condition requires $\nabla F(y_{K+1})^{\prime}(x-y_{K+1})\leq 0$ for all $x\in X_{T,K+1}$. Normalizing and using the fact that any $x\in X_{T,K+1}$ is a convex combination of points in $X_{A}$ and technology points $\{y_{1},\ldots,y_{K+1}\}$, the condition reduces to: $p_{100,K+1}^{\prime}y_{K+1}$ exceeds $p_{100,K+1}^{\prime}x$ for all extreme points $x$ of $X_{T,K+1}$. With strict inequality, $y_{K+1}$ is the unique optimum. $\square$
A.5  CES-CET closed forms
The closed-form expressions stated in Section 2 are derived here.
CES production function. The function $F(x;\theta,\sigma)=\left(\sum_{i}\theta_{i}^{1/\sigma}x_{i}^{(\sigma-1)/\sigma}\right)^{\sigma/(\sigma-1)}$ satisfies constant returns to scale, concavity for $\sigma>0$, and $C^{\infty}$ differentiability on $\mathbb{R}^{N}_{++}$.
CET resource use function. The function $g(x;s,\gamma)=\left(\sum_{i}s_{i}^{-1/\gamma}x_{i}^{(\gamma+1)/\gamma}\right)^{\gamma/(\gamma+1)}$ is convex, homogeneous of degree 1, and increasing. These properties follow from the convexity of $x^{(\gamma+1)/\gamma}$ for $\gamma>0$ and the composition rules for convex functions.
Autarky allocation. Setting up the Lagrangian $\mathcal{L}=F(x)-\mu(g(x)-B)$ and taking the ratio of first-order conditions for tasks $i$ and $j$:
$\frac{x_{i}}{x_{j}}=\left(\frac{\theta_{i}}{\theta_{j}}\right)^{\sigma/(\gamma+\sigma)}\left(\frac{s_{i}}{s_{j}}\right)^{\gamma/(\gamma+\sigma)}.$
Substituting into the constraint $g(x)=B$ yields the normalization $\Phi=\sum_{j}s_{j}^{(\sigma-1)/(\gamma+\sigma)}\theta_{j}^{(\gamma+1)/(\gamma+\sigma)}$, giving the closed-form allocation in the text.
Productivity index. By constant returns to scale, $F(x^{A})/B=F(x^{A}/B)$. Substituting the allocation formula and simplifying yields $F(x^{A})/B=\Phi^{\sigma/(\sigma-1)}$.
Shadow prices. From the gradient of $g$ at $x^{A}$: $p_{i}^{A}\propto s_{i}^{-1/\gamma}(x_{i}^{A})^{1/\gamma}$. Substituting the allocation formula and simplifying: $p_{i}^{A}\propto(\theta_{i}/s_{i})^{1/(\gamma+\sigma)}$, normalized to $\|p^{A}\|_{2}=1$.
Time shares. From the allocation formula, $\omega_{i}^{A}=x_{i}^{A}/\sum_{j}x_{j}^{A}\propto s_{i}^{\gamma/(\gamma+\sigma)}\theta_{i}^{\sigma/(\gamma+\sigma)}$. Logarithmic differentiation gives $\partial\omega_{i}^{A}/\partial s_{i}=[\gamma/((\gamma+\sigma)s_{i})]\omega_{i}^{A}(1-\omega_{i}^{A})>0$.
Unit revenue. The unit revenue function for CET is $\varrho(p)=(\sum_{i}s_{i}\,p_{i}^{\gamma+1})^{1/(\gamma+1)}$ (the dual $L^{q^{\prime}}$ norm with $q^{\prime}=\gamma+1$).
Entry threshold. From Definition 2.8, $\chi_{0}(t)=\underline{\chi}(t,p_{A})=\varrho(p_{A})/(p_{A}^{\prime}t)$. Substituting $p_{i}^{A}=\xi_{i}/\|\xi\|$ with $\xi_{i}=(\theta_{i}/s_{i})^{1/(\gamma+\sigma)}$, the $\|\xi\|$ terms cancel, giving $\chi_{0}(t)=\Phi^{1/(\gamma+1)}/\sum_{i}\xi_{i}t_{i}$ where $\Phi=\sum_{j}s_{j}^{(\sigma-1)/(\gamma+\sigma)}\theta_{j}^{(\gamma+1)/(\gamma+\sigma)}$.
Corner threshold. From Definition 2.8, $\chi_{100}(t)=\underline{\chi}(t,p_{100})$ where $p_{100}=\nabla F(t)/\|\nabla F(t)\|$. Under CES, $\nabla_{i}F(x)\propto\theta_{i}^{1/\sigma}x_{i}^{-1/\sigma}$, so $p_{100,i}\propto\theta_{i}^{1/\sigma}t_{i}^{-1/\sigma}$. Substituting into $\varrho(p_{100})$ and $p_{100}^{\prime}t$:
$\chi_{100}(t)=\frac{\left(\sum_{i}s_{i}\,\theta_{i}^{(\gamma+1)/\sigma}\,t_{i}^{-(\gamma+1)/\sigma}\right)^{1/(\gamma+1)}}{\sum_{i}\theta_{i}^{1/\sigma}\,t_{i}^{1-1/\sigma}}.$
The numerator is the worker’s unit revenue at technology-induced prices; the denominator is the technology’s value at those prices. Skills $s$ enter only the numerator: a more skilled worker has higher $\chi_{100}$ (harder to push to all-in).
Comparative statics. Both $\gamma$ and $\sigma$ enter $\chi_{0}$ through $\xi_{i}=(\theta_{i}/s_{i})^{1/(\gamma+\sigma)}$ and through the outer exponent $1/(\gamma+1)$. Higher $\gamma+\sigma$ compresses $\xi_{i}$ toward uniformity, making autarky prices less dispersed. The net effect on $\chi_{0}$ depends on the technology direction: technologies aimed at the worker’s weak tasks (high $\theta_{i}/s_{i}$) see the denominator $\sum_{i}\xi_{i}t_{i}$ fall, potentially raising $\chi_{0}$.
A.6  Proof of Corollary 3.1
(1) Cone width increases with $\gamma+\sigma$.
The cone of adopted directions at capability $\chi$ is $\mathcal{C}(\chi)=\{t\in\mathbb{S}^{N-1}_{+}:p_{A}^{\prime}t>\varrho(p_{A})/\chi\}$. The width of this cone depends on how concentrated $p_{A}$ is: a more uniform $p_{A}$ is positively aligned with more directions $t$.
Under CES-CET, $p_{i}^{A}\propto(\theta_{i}/s_{i})^{1/(\gamma+\sigma)}$. Define $\alpha=1/(\gamma+\sigma)$. As $\gamma+\sigma\to\infty$, $\alpha\to 0$, so $(\theta_{i}/s_{i})^{\alpha}\to 1$ for all $i$. Hence $p^{A}\to(1/\sqrt{N},\ldots,1/\sqrt{N})$—the uniform direction.
A uniform price vector satisfies $p_{A}^{\prime}t=1/\sqrt{N}\cdot\sum_{i}t_{i}>0$ for any $t\in\mathbb{S}^{N-1}_{+}$ with at least one positive component. Hence the cone converges to the entire positive orthant (a hemisphere in $\mathbb{S}^{N-1}$).
Conversely, as $\gamma+\sigma\to 0$, $\alpha\to\infty$, and $p^{A}$ concentrates on the task $i^{*}=\arg\max_{i}\theta_{i}/s_{i}$. Only technologies with large $t_{i^{*}}$ are adopted, making the cone very narrow.
(2) Partial-adoption region narrows with $\gamma+\sigma$.
By Lemma A.1, $\chi_{100}/\chi_{0}>1$ when $t\not\propto x_{A}$. The ratio measures how much the supporting hyperplane rotates from $x_{A}$ to $t$:
$\frac{\chi_{100}}{\chi_{0}}=\frac{\varrho(p_{100})/(p_{100}^{\prime}t)}{\varrho(p_{A})/(p_{A}^{\prime}t)}.$
As $\gamma+\sigma\to\infty$, both $p_{A}$ and $p_{100}$ converge to uniformity (the exponents $1/(\gamma+\sigma)$ and $1/\sigma$ in their definitions both shrink). When $p_{A}\approx p_{100}$, the thresholds $\chi_{0}$ and $\chi_{100}$ nearly coincide, so $\chi_{100}/\chi_{0}\to 1$.
In the rigid limit ($\gamma+\sigma$ small), $p_{A}$ and $p_{100}$ point in very different directions (one concentrated on the worker’s bottleneck, the other on the technology’s strength), widening the partial-adoption region.
Individual parameter effects. Holding other parameters fixed: (i) $\gamma\to 0$ (rigid specialist) allows a wide partial-adoption region; $\gamma\to\infty$ compresses it. (ii) $\sigma\to 0$ (Leontief production) allows a wide region; $\sigma\to\infty$ compresses it. Both effects work through the concentration of the price vectors $p_{A}$ and $p_{100}$. See Derivation M58 for the connection to the isoquant-scaling proof. $\square$

### v2

@@
 Abstract
-Generative AI is a directional technology: it excels at some task combinations and performs poorly at others.
-Knowledge work is also directional and endogenous: workers can satisfy their job requirements with different combinations of tasks.
-Studying AI adoption by knowledge workers hence requires comparing two vectors.
-We develop a high-dimensional model of task choice and technology adoption, with otherwise standard neoclassical assumptions.
-AI is adopted when its direction is aligned with what the worker values at the margin—the worker’s shadow prices, rather than with what the worker actually does—their activity vector.
-This yields a cone of adoption that widens as AI capability grows; near the entry threshold, small improvements in capability translate into large expansions in the set of adopted directions.
-Adoption also has a structured intensive margin: a tool can be worth using but not worth using all the time, generating a region of stable hybrid production between an entry threshold and an all-in threshold.
-We also show how to derive shadow prices as explicit functions of observable skill and requirement vectors.
-The framework explains rapid adoption in aligned occupations, heterogeneous adoption elsewhere, and weak correlation with one-dimensional skill measures: the key heterogeneity is directional alignment, not skill level.
+Generative AI is directional: it performs well in some task directions and poorly in others.
+Knowledge work is directional and endogenous as well: workers can satisfy the same job requirements with different mixes of tasks.
+We develop a high-dimensional model of AI adoption in which a worker uses a tool when it raises their output.
+Both the worker and the AI tool can perform a variety of tasks, which we model as convex production possibility sets.
+Because the tool requires supervision from the worker’s own time and attention budget, adoption is a team-production decision, similar to hiring a coworker.
+The key sufficient statistics are the worker’s pre-AI shadow prices: these equal the output gain from a small relaxation in each task direction, and they generally differ from the worker’s observed activity mix.
+As AI capability improves, the set of adopted directions expands in a cone centered on these autarky prices.
+Near the entry threshold, small capability improvements generate large extensive-margin expansions in adoption.
+The model also delivers a structured intensive margin: between the entry and all-in thresholds, optimal use is partial.
+We parametrize the model in a simple but flexible way that nests most existing task-based models of technical change.
 1  Introduction
-Generative artificial intelligence (GenAI) based on large language models (Vaswani et al., 2017) shares features of general-purpose technologies (Eloundou et al., 2023) but differs from earlier automation in two ways.
-First, individual workers can adopt GenAI quickly in knowledge work: the interface is general-purpose and the main inputs are time and attention rather than specialized capital.
-Adoption can therefore spread through informal and unregulated channels inside organizations.
+Generative artificial intelligence (GenAI) based on large language models (Vaswani et al., 2017) is not a single “machine.” It is a directional technology: it excels at some task combinations and performs poorly at others.
+Like other general-purpose technologies (Eloundou et al., 2023), it may reshape broad swaths of work, but two features make adoption and gains unusually hard to anticipate ex ante.
+First, adoption in knowledge work is often worker-driven: the interface is general-purpose and the main inputs are time and attention rather than specialized capital, so usage can spread inside organizations without explicit firm-level decisions.
 [^1] On worker-driven adoption (”shadow IT”), see Klotz et al.
@@
 (2025).
-Taken together, these features make it harder to predict who adopts AI and who gains from it.
-Measured productivity effects are mixed and do not map cleanly to any single notion of skill.
-In some settings, usage is already pervasive and mediated by tools that translate intent into working output; in others, adoption is slower and gains depend on complementary investments by firms.
-[^3] Evidence on productivity effects is mixed across settings and designs; see, e.g., Dell’Acqua et al.
-(2023); Benzell and Myers (2026) and firm-level evidence on complementarity in Aldasoro et al.
+Together, worker-driven adoption and jagged capability mean that asking “what is the effect of AI on labor?” is ill-posed without specifying which tools help with which tasks—and for which workers.
+Adoption is especially hard to characterize in knowledge work because workers can also change how they do the job.
+Job requirements may be stringent in terms of expected output, but different task mixes can deliver the same valued output.
+A software engineer can allocate time between writing new code, debugging, documentation, and code review; a management consultant between data analysis, client meetings, and slide preparation; a lawyer between legal research, drafting, and client communication.
+Reallocating effort across tasks is costly, however, and using AI consumes the same scarce time and attention that workers devote to their own work.
+To understand AI adoption incentives in knowledge work, we develop a model in which what a worker can do in a unit of time is a convex set of feasible task combinations—a production possibility set (PPS).
+[^3] This move is standard in international trade, where convex feasible sets and their supporting prices summarize production opportunities (Dixit and Norman, 1980).
+A worker is employed in a job, which determines the value of any bundle of tasks (Autor, 2013, 2025; Lise and Postel-Vinay, 2020).
+[^4] For evidence that workers have multidimensional skill endowments that affect occupational mobility, see Gathmann and Schoenberg (2010).
+They choose a task mix to maximize output subject to a time budget.
+Skills determine which task combinations are feasible; job requirements determine which combinations generate high output.
+We then treat AI as another production possibility set: a potential coworker with its own feasibility profile, and we model it as a directional technology that produces tasks in a fixed proportion.
+AI capability determines how much it can produce per unit of supervision time, in that fixed proportion.
+Using AI requires supervision, which uses time the worker could otherwise spend on their own tasks.
+The combined feasible set is therefore a weighted sum of what the worker and the tool can each produce, where the weights reflect how the worker allocates time between doing and supervising.
+Adopting AI is a collaboration decision rather than a simple substitution decision: the worker adopts when the expanded set is valuable enough to justify supervision.
+[^5] The same logic extends to tools that can flexibly combine a small number of task directions in a high-dimensional space; see Kaushik et al.
 (2025).
-Existing frameworks often compress heterogeneity into task or worker groups or a scalar exposure index, which can obscure the direction-by-direction nature of adoption and motivate a multidimensional model.
-To address this, we develop a model where both jobs and AI systems are explicitly multidimensional.
-A job is a bundle of tasks (Autor, 2013, 2025), and a worker chooses her task mix to maximize output subject to a time or attention budget.
-Skills shape the feasible set of task combinations, while job requirements shape the value of tasks in production.
-We model AI as a directional technology that produces tasks in fixed proportions.
-Using AI requires supervision—time the worker cannot spend on her own production.
-Because supervision draws on the same scarce budget, adoption is a team-production problem: the combined feasible set is a weighted sum of the worker’s own capabilities and those of AI.
-When their capabilities point in different directions, the joint production set expands beyond what either achieves in isolation.
-[^4] The logic extends to AI that can flexibly combine a small number of task directions in a high-dimensional space; see Kaushik et al.
-(2025).
-The key analytical step is to evaluate a technology direction at the worker’s marginal valuation: at the optimum, shadow prices convert any task direction into a payoff-relevant scalar.
-The model yields a simple organizing prediction: adoption is governed by alignment between the AI direction and what the worker values at the margin.
-AI is adopted when its direction is sufficiently aligned with the worker’s shadow prices, which are in general different from the worker’s observed activity mix.
-We make six contributions.
-  1. 1.
-    Explain nonlinear adoption without frictions.
-The model generates a sharp capability threshold for adoption even in a frictionless neoclassical setting.
-Near this threshold, small improvements in AI capability produce large expansions in the set of adopted tasks.
-This provides a clean economic explanation for why usage can suddenly take off without invoking hype, network effects, or fixed costs.
-  2. 2.
-    Separate absolute from comparative advantage.
-Technological superiority is necessary but not sufficient: AI is used only when its strengths align with what is scarce at the margin in a given worker–job match.
-This explains why powerful tools are sometimes ignored and why adoption varies widely across firms and roles.
-  3. 3.
-    Reframe AI as team production.
-Because AI requires supervision and consumes time and attention, it competes for the same scarce resource as human work.
-Adoption is therefore a collaboration decision evaluated at shadow prices, not a simple capital–labor substitution story.
-This shifts the discussion from replacement to directed co-production.
-  4. 4.
-    Introduce a structured intensive margin.
-There are two adoption thresholds in our model: an entry threshold and an all-in threshold.
-Between them, optimal AI use is partial, generating stable hybrid production rather than immediate automation.
-  5. 5.
-    Establish the worker–job match as the unit of analysis.
-Adoption depends on the interaction between worker skills and job requirements, not on workers or occupations in isolation.
-Shadow prices are match-specific, so asking which jobs or which workers are at risk is incomplete without specifying the pairing.
-  6. 6.
-    Break the scalar exposure paradigm.
-Adoption cannot be summarized by a one-dimensional skill or exposure index.
-The relevant sufficient statistic is a vector of shadow prices reflecting multidimensional alignment.
-As a result, common exposure rankings may systematically misclassify who adopts and who benefits.
-These results imply that scalar exposure rankings are generally not sufficient statistics for adoption.
-To make the mechanism transparent, our method is intentionally neoclassical.
+The worker adopts an AI tool if using it raises their output.
+At their pre-AI optimum, each task direction has a shadow (autarky) price: the output gain from being able to do slightly more of that task, holding everything else fixed.
+These shadow prices summarize what is scarce in the worker–job match and map the AI tool’s direction—a vector in task space—into a single scalar value.
+The worker adopts the AI tool when its direction is valuable at autarky prices.
+Autarky prices need not match the worker’s activity vector: they reflect what they would like to shift toward at the margin, not what they currently do.
+As AI capability improves, the set of adopted directions expands in a cone centered on the worker’s autarky prices; near the entry threshold, small capability gains translate into large extensive-margin expansions in adoption.
+Adoption also has a structured intensive margin: a tool can be worth using but not worth using all the time, generating a region of stable partial adoption between an entry threshold and an all-in threshold.
+Because autarky prices depend jointly on worker skills and job requirements, adoption is match-specific: neither “which workers adopt” nor “which jobs are affected” has a well-defined answer without specifying the pairing.
+Finally, greater flexibility—in how the worker can allocate their time or in how different tasks are valued in output—widens the set of tool directions that are worth adopting, so adoption expands faster as capabilities improve.
+We keep the analysis intentionally neoclassical to make the mechanism transparent.
 We study a worker doing a given job and abstract from governance and principal-agent frictions, so the worker maximizes output.
-We characterize autarky task choice using a trade-style duality argument (Dixit and Norman, 1980), which yields shadow (autarky) prices even though tasks are not traded on any market.
-We then introduce AI as a directional technology that competes for the same scarce budget and derive adoption and intensity conditions in terms of capability and alignment.
-For interpretation and measurement, we provide a CES-CET parametric specialization that expresses shadow prices as explicit functions of skill and requirement vectors.
-The key parameters are the elasticity of substitution across tasks in production ($\sigma$) and the elasticity of transformation in resource use ($\gamma$).
-When $\gamma\to 0$, workers are rigid specialists and skills alone determine autarky prices; when $\gamma\to\infty$, workers are perfectly flexible and job requirements alone matter.
-Symmetric logic applies to $\sigma$: low values make tasks strong complements, high values make them close substitutes.
-In general, both worker skills and job requirements shape autarky prices, consistent with within-job task reallocation (Spitz-Oener, 2006).
-The framework also has limitations that suggest where the economics of GenAI must ultimately go.
-By design, we hold the worker-AI match fixed and treat adoption as an individual optimization problem.
-We do not model sorting—which tools a worker selects and which workers a tool is designed for—or how organizations structure human-AI teams.
-In practice, firms set policies, provide complementary investments, and may restrict or encourage usage (Aldasoro et al., 2026).
-We also treat the technology as exogenous and focus on the worker side of adoption rather than on how training data and market incentives shape the direction of AI capabilities.
-These abstractions isolate a basic mechanism—directional alignment at shadow prices—that can be embedded in richer models of organizational choice, market structure, and dynamic capability change.
+The main objects and results—autarky prices, the adoption condition, the adoption cone, and the intensive margin—do not depend on a particular parametric form; they follow from convexity and marginal valuation.
+[^6] Formally, we characterize task choice using a trade-style duality argument that delivers supporting prices for a convex feasible set (Dixit and Norman, 1980).
+For measurement and interpretation, we also provide a parametric illustration using standard constant-elasticity functional forms.
+We assume a constant-elasticity-of-substitution structure for how tasks combine in output, and a constant-elasticity-of-transformation (CET) structure (Powell and Gruen, 1968) for how a worker’s fixed time/attention budget can be used for different task bundles.
+This specialization yields closed-form expressions for autarky prices as functions of observable skill and requirement vectors and clarifies when skills versus job requirements dominate adoption predictions.
+In the limiting cases, rigid specialists have shadow prices pinned down by job requirements, while perfectly flexible workers have shadow prices pinned down by skills.
+In between, which we believe is the relevant case for knowledge work, both matter.
+We hold the worker–job match fixed and analyze adoption of a given tool as an individual optimization problem.
+This isolates a building block that can be embedded in richer settings.
+In equilibrium, the availability of AI can reshape sorting and the distribution of jobs; multiple tools raise a team-composition problem; firms may govern usage and provide complementary investments (Aldasoro et al., 2026); and tool directions and capabilities may themselves respond to training data and market incentives.
 The remainder of the paper proceeds as follows.
-Section 2 develops the model of task choice and directional technology adoption, derives the adoption condition and the cone of adoption, and characterizes the intensive margin.
-Section 3 provides a CES-CET parametric specification that yields closed-form expressions for shadow prices and adoption thresholds.
-Section 4 positions the framework relative to prior task-based models and discusses interpretation, limitations, and connections to the emerging evidence.
-The Appendix collects proofs and the parametric derivations.
+Section 2 develops the high-dimensional model of task choice and directional technology adoption and derives the adoption condition, the adoption cone, and the intensive margin.
+Section 3 provides the parametric specification and its closed-form expressions.
+Section 4 relates the framework to prior task-based models and discusses interpretation and limitations.
+The Appendix collects proofs and additional derivations.
 2  A Model of Digital Technology Adoption
 We study a worker doing a given job.
-We do not model labor market, governance, or principal-agent frictions; consequently, the worker maximizes output.
-This is the outcome of a richer model in which both employer and worker can observe and contract on task execution.
-We take the worker-job match as fixed and reserve the study of sorting to future work.
+We abstract from labor-market, governance, and principal-agent frictions, so the worker maximizes output.
+This corresponds to a richer environment where tasks are observable and contractible within the firm.
+We take the worker–job match as fixed and reserve the study of sorting for future work.
 The worker chooses how much to do of each of $N$ tasks.
 Denote the task quantities by the vector $x\in\mathbb{R}_{+}^{N}$.
-The set of task bundles the worker can complete in one workday (or other unit of scarce resource) is given by the set $X_{A}$, characterized by its resource use function $g$,
+The set of task bundles the worker can complete in a workday (or other unit of the scarce resource) is given by the set $X_{A}$, characterized by its resource use function $g$,
 $X_{A}=\{x:g(x)\leq B\}.$
@@
 We interpret the budget as time or attention, not money.
-This is a key feature of digital technology adoption: the binding constraint is typically the worker’s cognitive capacity and hours, not financial cost.
+This is a key feature of digital technology adoption: the binding constraint is often the worker’s cognitive capacity and hours, not financial cost.
 The set $X_{A}$ is called the Production Possibilities Set (PPS).
@@
 (2026) leverage Becker’s time allocation theory to study consumer demand for time-intensive platforms like Facebook and Instagram, finding that time shares and time costs significantly influence substitution patterns across online activities.
-We assume AI is competing for the same resource budget $B$, so if the worker spends a fraction $\lambda$ of their budget working with AI, only $(1-\lambda)B$ is left for doing other tasks.
+We assume AI is competing for the same resource budget $B$, so if the worker spends a fraction $\lambda$ of their budget supervising AI, only $(1-\lambda)B$ is left for doing other tasks.
 [^9] For simplicity, we assume there is a linear transformation between resources spent on AI supervision and resources spent on human work.
@@
 In their current job, however, their activities are directionally far from the technology, so it is not worth adopting it.
-[^9] Geometrically, until the technology vector pierces the budget set given by the hyperplane separate the PPS and the production function upper contour set, its convex hull with the PPS is also strictly within the budget set, extending no revenue increasing opportunities to the worker.
+[^11] Geometrically, until the technology vector pierces the budget set defined by the hyperplane separating the PPS from the production function upper contour set, the convex hull of the technology point and the PPS remains strictly within the budget set and offers no revenue-increasing opportunities to the worker.
 […]
@@
 Proposition 2.7 (Unique optimal technology use).
-Suppose $F$ is strictly concave and $C^{2}$. Let $y=B\chi t$ and $X_{T}=\mathrm{conv}(X_{A}\cup\{y\})$. Then:
+Suppose $F$ is strictly quasi-concave and $C^{2}$. Let $y=B\chi t$ and $X_{T}=\mathrm{conv}(X_{A}\cup\{y\})$. Then:
   1. 1.
@@
 Proof.
-Uniqueness of $\tilde{x}_{*}$ follows from strict concavity of $F$ over the convex set $X_{T}$.
+Uniqueness of $\tilde{x}_{*}$ follows from strict quasi-concavity of $F$ over the convex set $X_{T}$.
 The decomposition follows from the convex hull representation (Proposition 2.3).
-Uniqueness of $\lambda_{*}$ follows from strict concavity of $f(\lambda)=\max_{x\in X_{A}}F((1-\lambda)x+\lambda y)$, and $x_{*}$ is then pinned down by $x_{*}=(\tilde{x}_{*}-\lambda_{*}y)/(1-\lambda_{*})$ when $\lambda_{*}<1$.
-Uniqueness of $p_{100}$ follows from $C^{2}$ differentiability of $F$.
+Uniqueness of $\lambda_{*}$ follows from strict concavity of $f(\lambda)=\max_{x\in X_{A}}F((1-\lambda)x+\lambda y)$ (which holds because $F$ is strictly quasi-concave and HD1), and $x_{*}$ is then pinned down by $x_{*}=(\tilde{x}_{*}-\lambda_{*}y)/(1-\lambda_{*})$ when $\lambda_{*}<1$.
+Uniqueness of $p_{*}$ follows from $C^{2}$ differentiability of $F$.
 See Appendix A.1.
@@
 Theorem 2.8 (Interior adoption).
-Suppose $F$ is strictly concave, $C^{2}$, and homogeneous of degree one. Let $\lambda_{*}(\chi)$ denote the unique optimal technology intensity from Proposition 2.7. Then:
+Suppose $F$ is strictly quasi-concave, $C^{2}$, and homogeneous of degree one. Let $\lambda_{*}(\chi)$ denote the unique optimal technology intensity from Proposition 2.7. Then:
   1. 1.
@@
 Cases 1–2 follow from the boundary conditions of the univariate function $f(\lambda)=F[(1-\lambda)x_{*}+\lambda y]$.
-For case 3, the strict inequality $\chi_{100}>\chi_{0}$ follows from strict concavity: the supporting hyperplanes at $x_{A}$ and at the technology direction $t$ are distinct whenever $t$ is not collinear with $x_{A}$.
+For case 3, the strict inequality $\chi_{100}>\chi_{0}$ follows from strict quasi-concavity: the supporting hyperplanes at $x_{A}$ and at the technology direction $t$ are distinct whenever $t$ is not collinear with $x_{A}$.
 See Appendix A.2.
@@
   3. 3.
-    All-in: The worker goes all-in on technology $K+1$ if $\chi_{K+1}>\varrho_{K}(p_{100,K+1})/p_{100,K+1}^{\prime}t_{K+1}$ , and $\chi_{K+1}/\chi_{k}>p_{100,K+1}^{\prime}t_{k}^{\prime}p_{100,K+1}^{\prime}t_{1}$ for all k=1,2,,Kk=1,2,...,K, where $p_{100,K+1}=\nabla F(t_{K+1})/\|\nabla F(t_{K+1})\|$.
+    All-in: The worker goes all-in on technology $K+1$ if $\chi_{K+1}>\varrho_{K}(p_{100,K+1})/p_{100,K+1}^{\prime}t_{K+1}$ and $\chi_{K+1}/\chi_{k}>(p_{100,K+1}^{\prime}t_{k})/(p_{100,K+1}^{\prime}t_{K+1})$ for all $k=1,\ldots,K$, where $p_{100,K+1}=\nabla F(t_{K+1})/\|\nabla F(t_{K+1})\|$.
 Proof.
@@
 $F$ satisfies constant returns, strict quasi-concavity, and $C^{2}$ differentiability on $\mathbb{R}^{N}_{++}$.
-Skills (CET). The resource use function is constant elasticity of transformation (CET):
+Skills (CET). The resource use function is constant elasticity of transformation (CET) (Powell and Gruen, 1968):
 $g(x)=\left(\sum_{i=1}^{N}s_{i}^{-1/\gamma}\,x_{i}^{(\gamma+1)/\gamma}\right)^{\gamma/(\gamma+1)},$
@@
 Then the autarky allocation takes the form
-$x_{i}^{A}=B\cdot\frac{s_{i}^{\gamma/(\gamma+\sigma)}\,\theta_{i}^{\sigma/(\gamma+\sigma)}}{\Phi^{\sigma/(\gamma+\sigma)}},\qquad i=1,\ldots,N,$
+$x_{i}^{A}=B\cdot\frac{s_{i}^{\sigma/(\gamma+\sigma)}\,\theta_{i}^{\gamma/(\gamma+\sigma)}}{\Phi^{\gamma/(\gamma+1)}},\qquad i=1,\ldots,N,$
 and autarky output per unit budget satisfies
@@
 The exponents make the role of curvature transparent.
-When $\sigma$ is high, tasks are easier to substitute in production, so the worker tilts more strongly toward high-requirement activities (higher weight on $\theta_{i}$).
-When $\gamma$ is high, it is easier to reallocate effort across tasks inside the constraint set, so differences in skills translate more strongly into differences in chosen activity levels.
+When $\gamma$ is high, the worker can freely reallocate across tasks, so they tilt toward what the job requires (higher weight on $\theta_{i}$).
+When $\sigma$ is high, tasks are easily substitutable in production and requirements matter less, so the worker leans toward tasks where they are skilled (higher weight on $s_{i}$).
 The productivity index $\Phi$ has a simple alignment interpretation: output is high when the worker is relatively skilled in the tasks that the job values (roughly, when $s_{i}$ is high where $\theta_{i}$ is high).
@@
 This also highlights that activities, in general, are poor measures of skills.
-## Applying the adoption theorems under CES–CET
+3.5  Applying the adoption theorems under CES–CET
 We now express the key statements about technology adoption in this parametric environment.
@@
 4  Discussion and Relation to the Literature
-Task-based labor-market models have a long history, and our framework builds on several strands of this literature.
-To clarify what is new, we organize prior work along five dimensions: (i) job heterogeneity (which task directions matter across jobs); (ii) worker task elasticity $\gamma$ (how flexibly workers reallocate across tasks); (iii) production elasticity $\sigma$ (how tasks combine in output); (iv) skill heterogeneity (how workers differ in task-specific productivity); and (v) technology directionality (whether technical change favors particular task directions).
-4.1  Organizing Framework
-Our model has four primitives: job requirements $\theta\in\mathbb{R}^{N}_{+}$, worker skills $s\in\mathbb{R}^{N}_{+}$, transformation elasticity $\gamma>0$ (the curvature of the worker’s task-allocation frontier), and substitution elasticity $\sigma>0$ (the curvature of the production aggregator).
-Technology enters as a direction $t\in S^{N-1}$ and capability $\chi>0$.
-This nests several important special cases.
-Table 1: Task-based models: key parameters
-Paper Jobs $\gamma$ $\sigma$ Workers Technology
-Welch (1969) NN-dim 0 $F$ NN-dim
-Katz and Murphy (1992) NN-dim 0 $F$ NN-dim general
-—application 2 types 0 $\sigma$ 2 types factor-augmenting
-Autor et al. (2003) 2 tasks $\infty$ 1 1-dim routine-biased
-Lazear (2009) 2 skills 0 $\infty$ 2-dim
-Acemoglu and Autor (2011) continuum $\infty$ 1 3 types task threshold
-Firpo et al. (2011) 5 tasks $\infty$ >1>1 5-dim 5-dim
-Cortes (2016) 3 tasks $\infty$ 0 routine-biased
-Fortin and Lemieux (2016) NN-dim 0 $\infty$ NN-dim
-Acemoglu and Restrepo (2018) continuum $\infty$ $\sigma$ 1 type rotating threshold
-Acemoglu and Restrepo (2020) continuum $\infty$ 1 1 type task threshold
-Bárány et al. (2020) NN-dim 0 $\infty$ NN-dim
-Böhm (2020) $K$ tasks logit $\infty$ routine-biased
-—application 3 tasks logit $\infty$ routine-biased
-Cavounidis and Lang (2020) NN-dim $\sigma$ $\infty$ NN-dim
-Cavounidis et al. (2024) NN-dim $\infty$ $\sigma$ NN-dim within-occ
-Autor and Thompson (2025) 1-dim $\infty$ 0 1-dim 1-dim
-Freund and Mann (2025) NN-dim $\infty$ 1 NN-dim directional
-Althoff and Reichardt (2026) NN-dim $\infty$ $\sigma$ directional
-Benzell and Myers (2026) NN-dim 0 $\sigma$ NN-dim 1-dim
-Gans and Goldfarb (2026) $n$ tasks $\infty$ 1 1-dim
-Bárány and Holzheu (2026) NN-dim 0 quadratic NN-dim
-This paper NN-dim $\gamma$ $\sigma$ NN-dim directional
-Note: “Jobs” and “Workers” refer to the dimensionality of heterogeneity in job requirements and worker skills, respectively.
-This is distinct from the number of tasks $N$: a model with many tasks can still be one-dimensional if tasks are ordered by a single attribute (e.g., routine content, expertise level).
-Many models take corner cases on at least one of the two curvature margins.
-It is common to assume either Leontief aggregation ($\sigma=0$) or perfect substitution ($\sigma=\infty$), and either rigid task endowments ($\gamma=0$) or frictionless reallocation across tasks ($\gamma=\infty$).
-Allowing both $\gamma$ and $\sigma$ to be finite matters because shadow prices and the direction of marginal value depend on both worker flexibility and production complementarity.
-This is the key step behind our alignment-based adoption condition and the structured intensive margin.
-4.2  Multidimensional Heterogeneity
-A key feature is that jobs and workers differ along multiple dimensions that cannot be collapsed to a single index.
-This rules out simple rankings of “high-skill” versus “low-skill” or “routine” versus “nonroutine” as sufficient statistics.
-Welch (1969) is among the first to decompose labor into multiple “labor inputs”—what we now call tasks.
-Workers have fixed endowments of these inputs ($\gamma=0$), and a general production function $F$ aggregates them.
-This early framework already allows for $N$-dimensional heterogeneity in both jobs and workers.
-Firpo et al.
-(2011) take a step toward higher dimensionality, using five O*NET task measures.
-Their reduced-form specification writes wages as an inner product of skills and returns—analogous to our $\langle s,\theta\rangle$ structure.
-Importantly, because task prices vary across occupations, tasks are implicitly not fully traded ($\gamma<\infty$).
-Their Oaxaca-Blinder decomposition captures changes in task content across occupations, which corresponds to “rotation” in our framework.
-Cortes (2016) studies job polarization with three task types and Leontief production ($\sigma=0$).
-Workers choose tasks freely ($\gamma=\infty$), and routine-biased technical change shifts employment across task categories.
-Böhm (2020) estimates task prices using a generalized Roy model where workers sort across $K$ tasks via discrete choice (effectively $\gamma=\text{logit}$).
-Tasks are traded at economy-wide prices ($\sigma=\infty$), and the focus is on recovering the price effects of routine-biased technical change.
-Two recent papers treat skills as genuinely multidimensional objects rather than as a scalar index.
-Bárány et al.
-(2020) develop and estimate a multi-sector general-equilibrium model of education and sector-occupation choices; jobs are bundles of tasks and workers sort on a multidimensional skill vector that evolves with schooling and experience (Abstract).
-Bárány and Holzheu (2026) define worker specialization as the distance between a worker’s skill set and those prevalent in the labor market and study its wage-unemployment trade-off in a random-search model with mismatch penalties and skill complementarity (Abstract).
-Relative to our framework, these papers emphasize schooling, sorting, and search frictions, whereas we focus on within-job task choice and the alignment-based adoption decision.
-Benzell and Myers (2026) is among the first to model within-job production with high-dimensional, non-traded tasks.
-Workers have skill endowments ($\gamma=0$), and tasks combine with CES elasticity $\sigma$.
-Automation supersedes one task direction.
-They emphasize that in high dimensions, workers cannot be ranked by a single index.
-Our model generalizes by allowing $\gamma>0$ (workers can reallocate across tasks) and directional technology.
-4.3  Within-Job Task Reallocation
-A second feature of our model is within-job task reallocation: workers shift effort across tasks while holding the job fixed.
-This margin is central to understanding how AI changes work without changing employment.
-Lazear (2009) introduces skills with job-specific weights—a direct antecedent of our $(\theta,s)$ formulation.
-In his baseline model, workers have endowments of two skills ($\gamma=0$), and firms combine them linearly with heterogeneous weights ($\sigma=\infty$).
-All skills are general, but firms value them differently, generating apparent firm-specific human capital.
-The model does not include technology, but the skill-weights structure is foundational.
-Fortin and Lemieux (2016) extend this skill-weights approach to study inequality, combining it with the task framework of Acemoglu and Autor (2011).
-Spitz-Oener (2006) provides early measurement evidence that task content changes within occupations over time, with computer use associated with shifts toward nonroutine cognitive tasks.
-Cavounidis and Lang (2020) microfound dynamic skill formation in a multidimensional setting.
-In their static model, workers choose task bundles with elasticity $\gamma=\sigma$ (in our notation) and output sums tasks linearly ($\sigma=\infty$), with job-specific task weights corresponding to our $\theta$.
-Cavounidis et al.
-(2024) apply this framework empirically, documenting within-occupation shifts in task content (what we call “rotation”) and connecting productivity effects to an envelope condition (our shadow prices).
-Autor and Thompson (2025) model expertise as a scalar that determines how task bundles translate into output.
-Technology compresses the expertise distribution by making low-expertise workers more productive.
-The task space is implicitly high-dimensional but collapsed to a single index.
-Our framework preserves the full dimensionality, showing that scalar indices can mask important heterogeneity in alignment.
-Freund and Mann (2025) develop a model where workers have multidimensional skill portfolios and jobs bundle multiple tasks, using 38 O*NET task clusters in their application.
-Tasks are freely tradable ($\gamma=\infty$) and combine with Cobb-Douglas technology ($\sigma=1$).
-They emphasize alignment between worker skills and job requirements and show that one-dimensional “exposure” measures can miss important heterogeneity.
-Their wage formula resembles our productivity expression, and their notion of skill-job “correlation” parallels our inner product structure.
-Our framework generalizes by allowing finite $\gamma$ (non-traded tasks) and finite $\sigma$ (task complementarity).
-Althoff and Reichardt (2026) develop a closely related framework where workers allocate time across $N$ tasks ($\gamma=\infty$) and output is CES with elasticity $\rho$.
-Technical change is task-specific and directional, allowing them to study how AI shifts comparative advantage across occupations.
-They extend the analysis to occupational choice, showing how task-specific technology shocks induce reallocation across jobs.
-Gans and Goldfarb (2026) study an O-ring production function where workers allocate time across $N$ tasks.
-Automation makes workers redundant in specific tasks, but reallocation occurs within the job.
-This is close to our setup with $\gamma=\infty$ and $\sigma=1$.
-The key difference is that their technology is one-dimensional (automation of an adjacent range of tasks), whereas ours is directional (technology produces a bundle of tasks in fixed proportions).
-Lin (2025) provides direct evidence for within-job task reallocation in response to AI changes.
-Using prompt-level data from Midjourney, she shows that most of the output shift after an upgrade comes from users changing their prompts rather than from the capability change itself.
-This is evidence of rotation in response to changing AI capabilities, exactly the intensive margin our model emphasizes.
-4.4  Other Task-Based Models
-Several other papers develop task-based frameworks that inform our approach but emphasize different margins.
-Katz and Murphy (1992) extend the Welch (1969) framework to study wage inequality: their general model allows $N$ labor inputs with a generic concave production function; their empirical application collapses to two skill types (college/non-college) with CES aggregation ($\sigma\approx 1.4$).
-They treat technical change as shifts in labor demand without specifying direction.
-The skill-biased technical change literature (Katz and Murphy, 1992; Acemoglu, 2002) partitions workers into broad skill groups and models technology as factor-augmenting.
-This approach captures between-group inequality but cannot address within-occupation heterogeneity in adoption or gains.
-The routine-biased framework (Autor et al., 2003; Acemoglu and Autor, 2011) adds a task dimension: technology substitutes for routine tasks while complementing non-routine tasks.
-Workers choose task bundles, but the task space is one-dimensional (routine vs.
-non-routine), and tasks are traded at economy-wide prices ($\gamma=\infty$).
-Acemoglu and Restrepo (2018, 2020) develop a workhorse model where a continuum of tasks is traded at economy-wide prices.
-Automation shifts the threshold between tasks performed by labor and capital.
-The production elasticity $\sigma$ governs macroeconomic implications.
-Our framework nests this as the limiting case $\gamma\to\infty$ (tasks fully tradable) with technology “rotating” the threshold across task indices.
-Lise and Postel-Vinay (2020) develop a sorting model with multidimensional skills.
-Workers and jobs match in a frictional labor market.
-The focus is on sorting and human capital accumulation rather than technology adoption.
-Agrawal et al.
-(2026) emphasize that AI may relax bottlenecks rather than automate tasks.
-This is complementary to our approach: in our model, AI that produces tasks where the worker is constrained (high shadow price) relaxes the binding constraint, which is the geometric interpretation of bottleneck relief.
-The contribution of our framework is to combine five features that prior models typically treat separately: (i) a high-dimensional task space; (ii) finite worker task elasticity $\gamma$ (non-traded tasks and worker-specific shadow prices); (iii) finite production elasticity $\sigma$ (task complementarity); (iv) multidimensional skill heterogeneity $s$; and (v) directional technology.
-The interaction of these features yields the adoption cone, the alignment condition, and the intensive margin.
+Task-based labor-market models have a long history.
+In contrast to much of this literature, and in line with international trade theory (Dixit and Norman, 1980; Helpman and Krugman, 1985), we model what workers can do as a convex production possibility set.
+This allows workers to flexibly adjust the tasks they supply in response not only to job requirements but also to technical change.
+Here we discuss how our approach differs from earlier models and how these differences affect our understanding of AI adoption.
+Our framework rests on two key assumptions: task bundling and worker flexibility.
+Task bundling refers to models where jobs require multiple tasks that are not tradable outside the specific worker-firm relationship.
+Foundational skill-bundle models (Welch, 1969; Katz and Murphy, 1992; Lazear, 2009; Autor and Thompson, 2025) treat workers as endowed with multiple skills that firms combine with heterogeneous weights.
+Sorting models (Gathmann and Schoenberg, 2010; Lise and Postel-Vinay, 2020; Bárány et al., 2020; Bárány and Holzheu, 2026) study how workers with multidimensional skills match to jobs with multidimensional requirements.
+Acemoglu and Autor (2011); Fortin and Lemieux (2016); Cavounidis and Lang (2020); Benzell and Myers (2026) use such task bundling models to study recent waves of automation.
+Worker flexibility refers to whether workers can reallocate effort across tasks within a job.
+Models with flexibility include the routine-biased literature (Autor et al., 2003; Acemoglu and Autor, 2011; Cortes, 2016; Böhm, 2020), the automation-threshold models (Acemoglu and Restrepo, 2018, 2020; Chen and Liang, 2026), expertise compression (Autor and Thompson, 2025), and most recent AI papers (Freund and Mann, 2025; Althoff and Reichardt, 2026; Gans and Goldfarb, 2026; Cavounidis et al., 2024; Cavounidis and Lang, 2020).
+Most models assume perfectly elastic reallocation across tasks, which, in our model, corresponds to $\gamma=\infty$.
+This assumption, when tasks are not bundled results in market task prices, which are neither worker- nor job-specific.
+If tasks are bundled, each worker’s shadow prices are pinned down by their relative skills rather than by match-specific marginal valuations, and thus are not job-specific.
+Very few models work with finite worker flexibility among a set of bundled tasks.
+Among the papers we survey, only Cavounidis and Lang (2020) allow workers to reallocate across tasks at a positive but finite cost, leading to the same CET production possibilities set as in our model.
+Their focus, however, is dynamic skill formation rather than technology adoption.
+Our paper gives a general characterization of the technology adoption problem for any degree of worker and job flexibility, thereby naturally embedding endowment models ($\gamma=0$), market-based task models ($\sigma=\infty$) and any intermediate case.
+To the best of our knowledge, we are also the first to model AI adoption as a team production problem that makes the combined feasible set convex.
 5  Conclusion
-GenAI is an unusually easy-to-adopt general-purpose technology, but its usefulness varies sharply across directions in a high-dimensional task space.
-This directional heterogeneity makes one-dimensional notions of exposure or skill unreliable guides to either adoption or productivity gains.
-We model a job as a neoclassical task allocation problem: workers choose task bundles subject to a convex resource constraint, and the optimal allocation is characterized by a vector of shadow (autarky) prices that summarize which tasks are scarce at the margin in the current worker-job match.
-We model AI as a directional technology with a capability level.
-Adoption then reduces to a value comparison at shadow prices.
-Three implications follow.
-First, absolute advantage is necessary but not sufficient: even a superior tool may not be adopted when its strengths point away from what is valuable at the margin.
-Second, adoption has a sharp entry threshold; near it, small capability gains can produce large expansions in the set of adopted directions.
-Third, adoption has a structured intensive margin: because entry and all-in thresholds generally differ, optimal use can remain partial over a range of capabilities, generating stable hybrid production.
-These mechanisms are match-specific because shadow prices depend jointly on worker skills and job requirements, and they extend naturally to settings with multiple tools, where adoption rotates shadow prices and can raise the bar for subsequent technologies.
+We model AI adoption as collaboration under a shared time and attention budget.
+A worker’s feasible task combinations form a convex production possibility set; an AI tool adds another such set with its own directional capabilities.
+Because using the tool requires supervision, the relevant object is the weighted sum of the worker and tool sets, with weights determined by time allocation.
+Adoption is therefore neither pure substitution nor pure augmentation: it is a team-production choice with an endogenous intensity.
+This framing sharpens what can (and cannot) be predicted without context.
+The right sufficient statistic is the worker’s autarky (shadow) prices at their pre-AI optimum: they summarize what is scarce at the margin in the worker–job match and translate the tool’s multidimensional direction into a scalar value.
+Adoption occurs when the tool’s direction is valuable at these prices.
+Because shadow prices are about marginal valuations, they need not coincide with observed task allocations, skills, or job requirements taken separately.
+Two implications follow for empirical work on GenAI in knowledge jobs.
+First, match-specificity is central: neither “which workers adopt” nor “which jobs are exposed” is well-defined in isolation because the same tool can be valuable in one match and irrelevant in another.
+Second, adoption has structured margins.
+As capabilities improve, adoption expands along an extensive margin (the set of adopted directions) and an intensive margin (partial versus all-in use), so small capability improvements near the threshold can generate large changes in usage while leaving other matches unchanged.
+These features help explain why scalar exposure indices can misclassify both adoption and gains when the technological frontier is jagged.
 Appendix A Proofs and Derivations
@@
 Autarky allocation. Setting up the Lagrangian $\mathcal{L}=F(x)-\mu(g(x)-B)$ and taking the ratio of first-order conditions for tasks $i$ and $j$:
-$\frac{x_{i}}{x_{j}}=\left(\frac{\theta_{i}}{\theta_{j}}\right)^{\sigma/(\gamma+\sigma)}\left(\frac{s_{i}}{s_{j}}\right)^{\gamma/(\gamma+\sigma)}.$
+$\frac{x_{i}}{x_{j}}=\left(\frac{\theta_{i}}{\theta_{j}}\right)^{\gamma/(\gamma+\sigma)}\left(\frac{s_{i}}{s_{j}}\right)^{\sigma/(\gamma+\sigma)}.$
 Substituting into the constraint $g(x)=B$ yields the normalization $\Phi=\sum_{j}s_{j}^{(\sigma-1)/(\gamma+\sigma)}\theta_{j}^{(\gamma+1)/(\gamma+\sigma)}$, giving the closed-form allocation in the text.
@@
 Time shares.
-From the allocation formula, $\omega_{i}^{A}=x_{i}^{A}/\sum_{j}x_{j}^{A}\propto s_{i}^{\gamma/(\gamma+\sigma)}\theta_{i}^{\sigma/(\gamma+\sigma)}$.
-Logarithmic differentiation gives $\partial\omega_{i}^{A}/\partial s_{i}=[\gamma/((\gamma+\sigma)s_{i})]\omega_{i}^{A}(1-\omega_{i}^{A})>0$.
+From the allocation formula, $\omega_{i}^{A}=x_{i}^{A}/\sum_{j}x_{j}^{A}\propto s_{i}^{\sigma/(\gamma+\sigma)}\theta_{i}^{\gamma/(\gamma+\sigma)}$.
+Logarithmic differentiation gives $\partial\omega_{i}^{A}/\partial s_{i}=[\sigma/((\gamma+\sigma)s_{i})]\omega_{i}^{A}(1-\omega_{i}^{A})>0$.
 Unit revenue. The unit revenue function for CET is $\varrho(p)=(\sum_{i}s_{i}\,p_{i}^{\gamma+1})^{1/(\gamma+1)}$ (the dual $L^{q^{\prime}}$ norm with $q^{\prime}=\gamma+1$).